### v1

# The Impact of Differential Feature Under-reporting on Algorithmic Fairness

Abstract.
Predictive risk models in the public sector are commonly developed using administrative data that is more complete for subpopulations that more greatly rely on public services. In the United States, for instance, information on health care utilization is routinely available to government agencies for individuals supported by Medicaid and Medicare, but not for the privately insured. Critiques of public sector algorithms have identified such differential feature under-reporting as a driver of disparities in algorithmic decision-making. Yet this form of data bias remains understudied from a technical viewpoint. While prior work has examined the fairness impacts of additive feature noise and features that are clearly marked as missing, the setting of data missingness absent indicators (i.e. differential feature under-reporting) has been lacking in research attention. In this work, we present an analytically tractable model of differential feature under-reporting which we then use to characterize the impact of this kind of data bias on algorithmic fairness. We demonstrate how standard missing data methods typically fail to mitigate bias in this setting, and propose a new set of methods specifically tailored to differential feature under-reporting. Our results show that, in real world data settings, under-reporting typically leads to increasing disparities. The proposed solution methods show success in mitigating increases in unfairness.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; June 03–06, 2024; Rio de Janeiro, Brazilbooktitle: FAccT ’24: ACM Conference on Fairness, Accountability, and Transparency, June 03–06, 2024, Rio de Janeiro, Brazilprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Regional and local governments around the world are using their increasingly digitized data systems to develop AI-driven decision-support technologies. The hope is that these tools improve decision quality, reduce inefficiencies, eliminate fraud, and improve outcomes for their citizens (Engin and Treleaven, 2019; Levy et al., 2021). Often, these technologies take the form of predictive risk models that are trained on administrative data to assess the likelihood that a case will go on to have poor outcomes. Such models have been developed and deployed in criminal justice (Barnes and Hyatt, 2012), child welfare (Vaithianathan et al., 2017), welfare fraud detection (Van Bekkum and Borgesius, 2021), federal tax audits (Houser and Sanders, 2016; Black et al., 2022), homelessness services (Kithulgoda et al., 2022), health care (McCarthy et al., 2015), and many other settings.
Predictive risk models in the public sector have come under criticism over concerns that they are trained on biased data (Mayson, 2019; Richardson et al., 2019; Chouldechova et al., 2018; Berk et al., 2018). In this paper, we consider a specific form of bias. We use the term ‘differential feature under-reporting’ to describe the phenomenon whereby administrative data records are more complete for individuals who have more greatly relied on public services. In the United States, for instance, administrative records often contain medical claims data for those who receive services through public insurance programs (Medicaid / Medicare), but lack information on physical, mental and behavioral healthcare utilization for the privately insured. A lack of recorded medical claims for an individual in this context is often indistinguishable from instances in which no medical claims have been made. Eubanks (2018) writes in her critique of the Allegheny Family Screening Tool (AFST) used in screening child maltreatment referrals,“by relying on data that is only collected on families using public resources, the AFST unfairly targets low-income families for child welfare scrutiny.” We provide a technical analysis of this problem. First, we introduce a statistical model of data collection with differential feature under-reporting. We then present theoretical results that characterize the impact of under-reporting on disparities in selection rates across groups. We describe why standard missing data methods generally fail to mitigate unfairness and, instead, propose new methods based on augmented loss estimation and optimal prediction imputation that are tailored to the under-reporting setting. Lastly, we present empirical results on semi-synthetic and real world data. Our experiments show that, in practice, under-reporting usually leads to increasing disparities and our proposed mitigation methods alleviate this increase.
$G$$Z$$Y$$\xi$$X$$\hat{Y}$
[…]
$\xi\text{: Publicly insured ($\xi=1$) or privately insured ($\xi=0$)}$
$Y\text{: Health risk}$
[…]
Figure 1. We study a prediction model on feature vectors with differential under-reporting $X$ where true outcomes $Y$ are a function of the latent ‘true’ features $Z$. Missingness $\xi$ is influenced by group membership $G$. We consider both cases in which feature distributions vary by group membership and cases with $G\perp Z$. In our setting, missingness indicators $\xi$ are unobserved and group membership $G$ is only used for model evaluation and not as a feature. The graph reflects the dependencies at prediction time.
2. Background and related work
Under-reporting vs. missingness
The problem of differential feature under-reporting is illustrated in Figure 1. An individual’s risk prediction $\hat{Y}$ is formed based on observed administrative data features $X$ which are a mismeasured version of a “true” latent feature vector $Z$. We assume that certain features in $Z$, such as demographic information, are correctly observed, whereas others, such as use of mental health services, are only correct for individuals who rely on publicly funded services. Problematically, we generally lack indicators on who is privately or publicly funded and for which services. For indicators and count features, e.g. the number of episodes in inpatient mental health treatment in the past year, the mismeasured feature will simply show the value 00 for individuals who received those services through privately funded mechanisms. This means that when we observe $X_{j}=0$, we do not know whether $Z_{j}=0$, or if $Z_{j}\neq 0$ and the feature has been mismeasured. In the graph, the unobserved missingness indicators are denoted by $\xi$. This distinguishes the under-reporting setting from standard missingness, wherein $\xi$ is assumed to be fully observed.
Missing data literature distinguishes three types of mechanisms: (1) Missing Completely At Random (MCAR) where missing values are independent of both observed and unobserved data, (2) Missing At Random (MAR) where missingness depends on observed variables, and (3) Missing Not At Random (MNAR) where missing values are influenced by systematic factors that are not recorded in the data (Rubin, 1976). In administrative data, records are generally more available for individuals who rely more greatly on public services which often correlates with demographic attributes like race or income levels. Since demographic variables are typically excluded from modeling, this implies an MNAR setting.
Under-reporting in real-world applications
The problem of feature under-reporting extends beyond the administrative data context. In health applications, Electronic Health Record (EHR) data is often under-reported at different levels for different population sub-groups (Wu et al., 2023; Rajkomar et al., 2018; Gianfrancesco et al., 2018; Groenwold and Dekkers, 2020; Cismondi et al., 2013). Socioeconomically disadvantaged patients may be missing more diagnostic tests due to limited health care access (Ahmad et al., 2019; Arpey et al., 2017). Reliance on clinical decision support systems trained on EHR data could exacerbate already existing health care disparities (Gianfrancesco et al., 2018; Char et al., 2018; Jeanselme et al., 2022). Similarly, the extent of under-reporting often varies across domains (e.g. hospitals) which has been studied by Zhou et al. (2022). While Zhou et al. (2022) consider model adaptation when shifting to unlabeled target data with different level of under-reporting, we focus on a single domain with varying levels of under-reporting across groups and study fairness implications.
Under-reporting in biomedical research
In epidemiological surveys, social stigma can lead participants to provide false negative responses (e.g. true $Z$ vs. reported maternal smoking $X$) (King et al., 2001; McKnight et al., 2007; Greenland, 2014; Sechidis et al., 2017). Researchers have proposed various methods to estimate association between $Z$ and outcome $Y$ by leveraging $X$ including correction factors for independence tests (Sechidis et al., 2017; Bross, 1954), adjusted mutual information (Sechidis et al., 2017), odds-ratios (Chu et al., 2006; Edwards et al., 2013; Dosemeci et al., 1990), risk-ratios (Rahardja and Young, 2021; Brenner and Loomis, 1994), and full likelihood approaches (Adams et al., 2019). These methods are typically limited to binary features and outcomes, and focus on statistical inference rather than prediction. Even in settings in which a correctly specified model $f(z)=\mathds{E}\left[{Y\mid Z=z}\right]$ can be estimated using only $X$ and $Y$, it is unclear how to use this model when only under-reported features are available at prediction time. In the study of geographical disease counts, ‘under-reporting’ typically refers to cases with $X\leq Z$. Most work in this area employs Bayesian modeling (Bailey et al., 2005; de Oliveira et al., 2017; Stoner et al., 2019; Gelman et al., 2013) which requires a host of parametric and distributional assumptions as well as informative priors. In single-cell RNA sequencing, ‘zero-inflation’ refers to the phenomenon by which genes are undetected despite being expressed in a cell due to low levels of RNA. Correction methods typically employ Bayesian methods that are highly specialized for the single-cell RNA sequencing task.
Additive noise and fairness
Complete data Additive noise Missing with indicator Under-reporting
Setting
$g$ $\bm{z_{1}}$ $z_{2}$ $y$
0 10 2 1
0 7 1 0
1 0 3 1
$g$ $\bm{x_{1}}$ $z_{2}$ $y$
0 10.2 2 1
0 6.5 1 0
1 0.8 3 1
$g$ $\bm{x_{1}}$ $\bm{r}$ $z_{2}$ $y$
0 10 1 2 1
0 $\bm{m}$ 0 1 0
1 $\bm{m}$ 0 3 1
$g$ $\bm{x_{1}}$ $z_{2}$ $y$
0 10 2 1
0 $\bm{m}$ 1 0
1 $\bm{m}$ 3 1
Features fully observed
Feature values with added random noise $\varepsilon$
Some feature values take default value $m$; $r$ indicates which values are observed
Some feature values take default value $m$; No indicators for missingness
Previous fairness work
No feature mismeasurement
Khani and Liang (2020), Phelps (1972), Aigner and Cain (1977), Chen et al. (2018)
Zhang and Long (2021), Wang and Singh (2021), Jeanselme et al. (2022), Fernando et al. (2021), Fricke (2020), Ahmad et al. (2019)
This work, Eubanks (2018)
Table 1. Different types of feature mismeasurement and previous work addressing fairness implications. In the data examples, mismeasured features are denoted with $x$ while correctly observed features are denoted by $z$. Column $g$ encodes group membership.
Algorithmic fairness literature has studied various types of feature mismeasurement as summarized in Table 1. A commonly studied setting is additive feature noise where, instead of a feature $z_{1}$, we observe a noisy version $x_{1}=z_{1}+\varepsilon$. The random noise $\varepsilon$ is often assumed to be zero-mean, of small variance, and independent of other variables which implies that, while some of the information in the feature is diluted, large portions of the encoded information remains intact. Khani and Liang (2020) show that adding the same amount of feature noise to a group-blind model can introduce statistical loss discrepancy. This is in line with earlier observations from the statistical discrimination literature (Phelps, 1972; Aigner and Cain, 1977). Chen et al. (2018) propose data collection strategies targeted at decreasing discrepancy and come to the conclusion that overcoming differential noise across protected groups may require collection of additional data. In contrast to additive noise, under-reporting removes all information from impacted feature entries and typically biases the feature mean. Some works (e.g. Khani and Liang, 2020) suggest that feature missingness can be modeled as a special case of additive noise by selecting noise terms with very high variance. However, this only covers a special case of under-reporting in which feature entries are missing for all observations.
Missing data methods and fairness
Feature missingness has been studied in the statistical literature for several decades (e.g. Rubin, 1976). This line of work generally assumes that we observe an indicator variable for missing entries or, equivalently, missing values are clearly marked as such (e.g. as NaN). Various methods with different fairness implications have been proposed to overcome this types of missingness (see also Table 1).
  1. (1)
    Complete case analysis and reweighing. In some cases, it may be desirable to remove incomplete rows which can lead to significant biases (Rubin, 1976). Various reweighing procedures have been proposed to deal with this problem. Zhang and Long (2021) suggest learning only from complete observations while employing an importance sampling procedure. Wang and Singh (2021) suggest that reweighing and resampling methods in the context of categorical data can lead to considerable fairness improvements over learning with missing data directly.
  2. (2)
    Imputation. Jeanselme et al. (2022) compare different imputation strategies under clinical presence and find that there is no imputation strategy that reliably outperforms other imputation methods in terms of fairness. Fernando et al. (2021) and Fricke (2020) compare feature imputation to complete case analysis and find that rows with missing values can contribute to fairer outcomes via observed columns. Ahmad et al. (2019) raise concerns about potential harms of imputation in high-stakes settings like recidivism prediction.
In contrast to the feature missingness setting, we do not observe indicators for missing entries in this work. Despite this difficulty, we experiment with row omission and imputation in Section 7.
3. Problem setup
Setting
We study the effect of feature under-reporting on algorithmic fairness through the lens of the regression setting displayed in Figure 1. Assume latent feature vectors $z\in\mathds{R}^{d}$ and group information $g\in\{0,1\}$. We assume a noiseless regression setting in which the outcome $y$ is a linear function of $z$, i.e. $y=\alpha+\beta^{T}z$ with $\alpha\in\mathds{R}$, $\beta\in\mathds{R}_{\neq 0}^{d}$. Instead of the true features $z$, we observe a mismeasured vector $x$ in which entries default to 0 with group-dependent probabilities. That is, we set $x=z\odot\xi^{g}$ where $\odot$ denotes element-wise multiplication, $\xi^{g}\sim\text{Bern}(m^{g})$, and $(1-m^{0}),(1-m^{1})\in(0,1]^{d}$ are under-reporting rates in the two groups. More formally, we have a group random variable $G\sim\text{Bern}(r)$ and a random feature vector $Z$. The vector of under-reported features $X$ can be written as $X=Z\odot\xi$ where $\xi=G\xi^{1}+(1-G)\xi^{0}$. This setting allows for different dependence structures depending on whether we assume $G\perp Z$.
Two-step bias
Differential under-reporting introduces bias in two ways: (1) Under-reporting in training data influences the estimation of the prediction model (estimation step), and (2) input data with under-reporting leads to biased predictions at test and deployment time (prediction step). We note that, in general, it is not sufficient to recover the true model parameters $\alpha,\beta$ as only biased features are available at prediction time. In fact, our experiments in Section 8 demonstrate that using the true model parameters for prediction with under-reporting can lead to worse fairness outcomes than using a model estimated with biased data directly.
Thresholded prediction
We assume a thresholded prediction setting reminiscent of predictive risk modeling in the public sector. A predictor $f$ is fit on $(X,Y)$ to produce predictions $\hat{Y}=f(X)=f(Z\odot\xi^{G})$. We consider the group-wise shares of predictions that lie above a given threshold $\tilde{y}$: $\mathbb{P}(\hat{Y}\geq\tilde{y}\mid G=g)$ which we refer to as selection rates at threshold $\tilde{y}$. This setup implicitly assumes a predictive risk assessment setting in which the highest risk individuals are selected (e.g. child welfare screenings, fraud detection, federal tax audits). However, it is straightforward to reverse the analysis for scenarios in which low risk leads to selection (e.g. bail decisions in criminal risk assessment). In addition, we assume that being selected is undesirable and the group that is over-selected is disadvantaged. Crucially, this assumption is only made to simplify interpretation of results and we could easily consider the opposite case.
Excess selection rates
We select the top $C\in[0,1]$ share of the population by thresholding predictions $\hat{Y}$. Given the cumulative distribution function of predictions $F_{\hat{Y}}$, the percentile threshold $C$ implies an absolute threshold $\tilde{y}=F_{\hat{Y}}^{-1}(1-C)$ such that the selection rate for group $g$ can be written as $P(\hat{Y}\geq\tilde{y}\mid G=g)$. In order to isolate the effect of under-reporting, we need to account for a ground truth difference in selection rates. Let $\hat{Y}_{X}$ denote the predictions of a model trained on $(X,Y)$ and $\hat{Y}_{Z}$ the predictions of a model trained on $(Z,Y)$. Since the distributions of $\hat{Y}_{Z}$ and $\hat{Y}_{X}$ generally differ, the predictions imply different thresholds $y^{\prime}=F_{\hat{Y}_{Z}}^{-1}(1-C)$ and $\tilde{y}=F_{\hat{Y}_{X}}^{-1}(1-C)$. With this notation in hand, we are able to define an equity-related metric that allows us to assess the impact of differential feature under-reporting on disparities in selection rates.
Definition 1 (Excess selection rate due to under-reporting).
The excess selection rate for group $g\in\{0,1\}$ at overall selection rate $C\in[0,1]$
$\Delta(g,C):=P(\hat{Y}_{X}\geq\tilde{y}\mid G=g)-P(\hat{Y}_{Z}\geq y^{\prime}\mid G=g),$
is the difference in selection rates when ranking according to a model trained on $X$ compared to a model trained on $Z$. We say that group $g$ is over-selected at level $C$ if $\Delta(g,C)>0$. If $\Delta(g,C)<0$, we say that $g$ is under-selected.
In principle, we could directly consider a “difference in difference”: the difference in selection rates between groups $g=0,1$ when selection occurs according to the model $\hat{Y}_{X}$ versus the unbiased predictions $\hat{Y}_{Z}$. However, since we select a fixed share of the population $C$, an increase of the selection rate of group $g$ when moving from $\hat{Y}_{Z}$ to $\hat{Y}_{X}$ already implies a decrease for group $1-g$. It is generally difficult to argue about the excess selection rate $\Delta(g,C)$ analytically. Even in a simple setting with group-dependent Gaussian features $Z\mid G\sim\mathcal{N}\left(\mu_{G},\Sigma_{G}\right)$, there is no closed-form expression for the quantile $y^{\prime}=F_{\hat{Y}_{Z}}^{-1}(1-C)$ and determining the sign of $\Delta(g,C)$ requires analysis of a difference in cdfs which is often intractable. Instead, we simplify the setting and assume that $Z$ follows the same distribution across groups. In this case, the selection rates on the true outcome $Y$ are the same in both groups at every threshold, and we can simplify Definition 1 as follows.
Definition 2 (Excess selection rate due to under-reporting, independent case).
If $Z\perp G$, we say that group $g\in\{0,1\}$ is over-selected at threshold $C\in[0,1]$ if $P(\hat{Y}_{X}\geq\tilde{y}\mid G=g)>P(\hat{Y}_{X}\geq\tilde{y}\mid G=1-g),$ and under-selected if the inequality is reversed.
While the majority of our theoretical derivations assume the special case of $Z\perp G$, the empirical portion of this work explores the impact of feature under-reporting on selection rates in the more general setting.
4. Differential feature under-reporting in linear regression
In this section, we examine the bias that differential feature under-reporting introduces into parameter estimates in linear regression. We consider a setting in which true outcomes are a linear function of the latent feature vectors, i.e. $Y=\alpha+\beta^{T}Z$, which implies that a linear model with access to the true $Z$ recovers the true outcomes $Y$. To simplify notation, we drop the subscripts and write $Y=\hat{Y}_{Z}$ and $\hat{Y}=\hat{Y}_{X}$ for the remainder of the paper. Note that the theoretical portion of this paper considers population-level regressions and we do not discuss finite data results.
Estimates and attentuation bias
Feature mismeasurement in the form of under-reporting leads to inconsistent parameter estimates in linear regression. When fitting a linear model on $(X,Y)$, the least squares estimates become
(1) $\hat{\beta}$ $=\Sigma_{X}^{-1}\Sigma_{XZ}\beta,$ $\quad\hat{\alpha}$ $=\alpha+\mathds{E}\left[{Z}\right]^{T}\beta-\mathds{E}\left[{X}\right]^{T}\hat{\beta},$
where $\Sigma_{XZ}$ denotes the covariance matrix between $X$ and $Z$ and we write $\Sigma_{X}$ for $\Sigma_{XX}$. At first glace, this solution resembles the regression estimates in the more commonly studied additive feature noise case. Assuming $X^{\prime}=Z+U$ where $U$ is independent zero-mean feature noise, we receive $\hat{\beta}=\Sigma_{X^{\prime}}^{-1}\Sigma_{X^{\prime}Z}\beta=(\Sigma_{Z}+\Sigma_{U})^{-1}\Sigma_{Z}\beta$. The factor $\lambda=(\Sigma_{Z}+\Sigma_{U})^{-1}\Sigma_{Z}$ is commonly interpreted as a noise-to-signal ratio and, if $Z$ is one-dimensional, we know that $\mid\hat{\beta}\mid=\lambda\mid\beta\mid<\mid\beta\mid$ which is generally refered to as attentuation bias (e.g. Hausman, 2001; Fuller, 1987). In the under-reporting setting, $\Sigma_{X}$ does not easily separate into terms depending on only the feature or only the mismeasurement. However, in the special case of one-dimensional $Z$ and $Z\perp G$, we can still show that the parameter $\hat{\beta}$ is biased towards zero.
Lemma 1 (Attentuation bias).
Assume the feature $Z$ is one-dimensional and has the same distribution across groups. Then, the least squares regression of $Y$ on the mismeasured feature $X$ yields an estimated slope $\hat{\beta}$ with $\mid\hat{\beta}\mid\leq\mid\beta\mid$.
The $d$-dimensional case
Real-world prediction settings typically include several features that tend to be correlated in some manner. Assume the feature vector $Z$ is $d$-dimensional, and under-reporting only occurs in the first feature. This means that $X$ coincides with $Z$ in all but the first entry which is computed as $X_{1}=Z_{1}\xi_{1}$ where $\xi_{1}=G\xi_{1}^{1}+(1-G)\xi_{1}^{0}$ and $\xi_{1}^{1}\sim\text{Bern}(m^{0}_{1})$, $\xi_{1}^{0}\sim\text{Bern}(m^{1}_{1})$. We further assume that features $Z_{2},\ldots,Z_{d}$ are uncorrelated and the feature dependence structure is characterized entirely by the correlations between $Z_{1}$ and $Z_{[2:d]}$. In practical settings, this may be achieved via feature orthogonalization. We explicitly exclude cases in which $Z_{1}$ is a perfect linear combination of other features to avoid problems of multicollinearity and assume $\mathds{V}\left[{Z_{k}}\right]>0$ for all $k\in[2:d]$. Given this setting, Proposition A.1 provides a closed form representation for the parameter estimates $\hat{\beta}$. The bias in these estimates can be conceptualized as a generalization of omitted variable bias (Angrist and Pischke, 2008) which is discussed Appendix B. We now assume that $Z\perp G$ which implies that under-reporting is independent of the value that is under-reported. This resembles the assumptions made in previous work on the impact of additive feature noise on fairness (e.g. Khani and Liang, 2020; Phelps, 1972; Aigner and Cain, 1977), and allows us to gain analytical insights that would otherwise remain intractable. First, we examine the behavior of the parameter estimate for the feature with under-reporting, i.e. $\hat{\beta}_{1}$.
Proposition 2 (Properties of $\hat{\beta}_{1}$).
If $Z\perp G$, the parameter estimate $\hat{\beta}_{1}$ has the following properties.
  1. (1)
    Sign invariance: $\hat{\beta}_{1}$ has the same sign as $\beta_{1}$.
  2. (2)
    Attentuation bias: $\mid\hat{\beta}_{1}\mid\leq\mid\beta_{1}\mid$.
  3. (3)
    Attentuation bias increasing with under-reporting: If under-reporting $1-m_{1}^{g}$ is increasing for one (or both) groups $g\in\{0,1\}$, ceteris paribus, the magnitude of the parameter estimate $\mid\hat{\beta}_{1}\mid$ is decreasing.
This finding shows that feature under-reporting leads to attenuation bias in the respective parameter estimate even when other correlated and fully observed features are available. The attenuation bias gets more pronounced with more under-reporting. Next, we turn towards the estimates for the fully observed features $\hat{\beta}_{k}$ for $k\in[2:k]$.
Proposition 3 (Properties of $\hat{\beta}_{k}$).
If $Z\perp G$, the parameter estimates $\hat{\beta}_{k}$ for $k\in[2:d]$ have the following properties.
  1. (1)
    Correlation bias: If $\hat{\beta}_{k}\neq\beta_{k}$, then $\rho(Z_{1},Z_{k})>0$.
  2. (2)
[…]
In line with general intuition, under-reporting in $Z_{1}$ has no effect on the parameter estimate $\hat{\beta}_{k}$ if features $Z_{k}$ and $Z_{1}$ are uncorrelated. If the features are correlated, the direction of the under-reporting effect on the parameter estimate depends on the signs of $\beta_{1}$ and $\text{Cov}\left[{Z_{1}},{Z_{k}}\right]$. Note that this is independent of the value and sign of $\beta_{k}$.
Take-away
Proposition 2 and 3 tell a compelling story about the effect of under-reporting on parameter estimates in the studied setting. As more feature values default, the regression model places less weight on the mismeasured feature and instead shifts weight to fully observed features with non-zero correlation. This can lead to increasing or decreasing parameter estimates. Analytically, there are cases in which ‘shifting weight’ means that the absolute value of a parameter estimate $\mid\hat{\beta}_{k}\mid$ is decreasing which may appear counterintuitive. For example, consider a setting in which both $Z_{1}$ and $Z_{k}$ have positive true parameters but negative correlation, i.e. $\beta_{1},\beta_{k}>0$ and $\text{Cov}\left[{Z_{1}},{Z_{k}}\right]<0$. In practice, this could occur when there are several mutually exclusive paths to the same outcome. For example, consider prediction of general health risk scores with features including both the number of pediatrician visits in the last year and the number of internist visits. Presumably, these features are negatively correlated because they are relevant for two mutually exclusive parts of the population, i.e. children and adults, but in both columns larger values can be indicative of a high general health risk.
5. Impact on selection rate disparity
Selection rate disparity in Gaussian setting
We study the effect of differential feature under-reporting on selection rate disparities in linear regression. For the remainder of this section, we assume that $Z\perp G$. Similar to our previous discussion, we assume a $d$-dimensional feature setting in which only the first feature is subject to under-reporting while all other features are fully observed. We further assume that features are jointly Gaussian, i.e. $Z\sim\mathcal{N}\left(\mu,\Sigma\right)$ where $\mu\in\mathds{R}^{d}$ and the covariance matrix $\Sigma\in\mathds{R}^{d\times d}$ is positive definite. This has the benefit that predictions $\hat{Y}=\hat{Y}_{X}$ follow a Gaussian mixture distribution which allows us to directly analyze group selection rates. If under-reporting rates are the same across groups, i.e. $m_{1}^{0}=m_{1}^{1}$, there is no selection rate disparity as both groups have the same distributions of features and predictions. If the under-reporting rates vary between groups, we observe the following.
Proposition 1 ().
For a sufficiently high threshold $\tilde{y}$, the group with more under-reporting is over-selected if
$\mathds{V}\left[{\hat{\alpha}+\hat{\beta}_{[2:d]}Z_{[2:d]}}\right]>\mathds{V}\left[{\hat{\alpha}+\hat{\beta}Z}\right]$
(Case 1), or under-selected if the inequality is reversed (Case 2). For low thresholds, the cases are reversed.
We refer to Proposition A.2 for an expanded version of this finding including a discussion of sufficiently high thresholds. Proposition 1 shows that over-selection primarily depends on the variance in predictions. When cutting off at a high threshold, the group with more feature under-reporting is over-selected if the variance in predictions for examples with defaulted feature exceeds the prediction variance for fully observed examples (Case 1). It is under-selected if the variance of predictions is larger for the examples with fully observed features (Case 2). Intuitively, we may expect that, in prediction settings with high thresholds, information deficiency in a group always leads to under-selection of that group because it prompts the group’s risk distribution to concentrate more closely around its mean moving more mass below the threshold. We find that, analytically, outcome disparities can go into either direction and sometimes groups with information deficiency are over-selected. While our findings suggest that this is mostly a question of variance in predictions, this is likely only part of the story in more general cases with group-dependent feature distributions. More general settings are studied empirically in Section 7.
Combining parameter estimation and prediction steps
Feature under-reporting introduces bias both at estimation and prediction time. In the following, we combine our previous findings to examine the conditions under which under-reporting leads to over-selection and under-selection. As before, we assume a $d$-dimensional feature setting in which only the first feature $Z_{1}$ is impacted by under-reporting. Features are jointly Gaussian, and we further assume that $Z_{2},\ldots,Z_{d}$ are uncorrelated.
Corollary 2 ().
Given the first and second moments of $Z_{1}$, the expected share of observed values $\mathds{E}\left[{\xi_{1}}\right]$, and the fraction of variance in $Z_{1}$ that is explained by the remaining features $S^{2}=\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}$, there exists a positive constant $c=c(\mathds{E}\left[{Z_{1}}\right],\mathds{V}\left[{Z_{1}}\right],\mathds{E}\left[{\xi_{1}}\right],S^{2})$ such that, at high thresholds, the group with more under-reporting is over-selected if
[…]
(Case 1), and under-selected if the inequality is reversed (Case 2).
Thresholds are considered high if they exceed the turning point defined in Proposition A.2. The corollary shows that over-selection due to feature under-reporting depends on the signs and magnitudes of the true parameters $\beta$ and the covariances between features. If $\text{sign}\left(\frac{1}{\beta_{1}}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]\right)=1,$ If the sign is negative, the group with more under-reporting is over-selected if the covariance-weighted sum of true parameters is sufficiently large in absolute value. Otherwise, feature under-reporting still leads to under-selection.
6. Solution approaches
Sections 4 and 5 show how ignoring differential feature under-reporting can lead to unfair selection rates across groups. In the following, we explore how conventional missing data methods can be adapted to the under-reporting setting. We then propose a new set of methods that is specifically tailored to this setting by separating the problem into two steps — estimation and prediction. For the estimation step, we provide a method that recovers the ground truth data generating model from observed data. For the prediction step, we derive optimal group-dependent imputation values. As before, we assume under-reporting occurs only in the first feature which is observed as $X_{1}=Z_{1}\xi_{1}$.
Standard missing data methods
Existing missing data methods typically assume that defaulted values are clearly marked which is not the case in the under-reporting setting. We explore adaptations of several methods.
  1. (1)
    Feature omission. Discarding mismeasured feature vector $X_{1}$ entirely mitigates the bias introduced by under-reporting without requiring missingness indicators. While this may be a viable option in some settings, it can lower model performance significantly and the fairness impact is non-obvious. When assuming a linear ground truth, feature omission leads to omitted variable bias in parameter estimates $\hat{\beta}_{[2:d]}$ which has been studied previously in the econometrics literature (Angrist and Pischke, 2008) (also see Appendix B).
  2. (2)
    Multiple imputation. Multiple imputation draws plausible feature values while retaining variability. Since we do not observe indicators for missingness, we experiment with imputing all 0-entries in $X_{1}$ which includes correctly observed 0s. In each imputation run, we estimate the posterior $P(Z_{1}\mid Z_{2},\ldots,Z_{d})$ on data rows with $X_{1}\neq 0$, impute 0-entries, and train a prediction model for $Y$. At prediction time, we average the imputations and predictions over models to obtain a single prediction $\hat{Y}$. While this procedure successfully alleviates bias in some standard feature missingness settings, it is not a priori clear how well the method works with under-reporting.
  3. (3)
    Row omission. Omission of rows with missing feature entries provides a convenient complete case analysis. Since true and false 0-entries are indistinguishable in our setting, we experiment with discarding all rows with $X_{1}=0$. If there is no model misspecification, e.g. in the linear case with $f(Z)=\alpha+\beta^{T}Z$ where we train a linear model on observed features, training on only complete rows is guaranteed to asymptotically retrieve the true parameters $\hat{\alpha}=\alpha$ and $\hat{\beta}=\beta$ if $Z_{1}$ is not binary. Even with access to the ground truth model, under-reporting introduces bias via the prediction step and $\hat{Y}=f(X)$ may not be the most accurate (or fairest) prediction.
Model estimation with augmented loss
Without model misspecification, row omission can recover the ground truth model $Y=f(Z)$. In practice, models are usually misspecified and discarding rows can significantly decrease performance. Instead, we propose an augmented loss function to recover the ground truth model. This proxy loss uses observed features $X$ to provide an unbiased estimate of the loss of a model $f$ on latent features $Z$. Similar approaches have previously been used in the label noise setting (Natarajan et al., 2013; Nemirovski et al., 2009). Assume $Z\in\mathds{R}^{d}$ has support $\mathcal{Z}$ and $y\in\mathds{R}$ has support $\mathcal{Y}$. Let $\mathcal{F}:\mathcal{Z}\to\mathds{R}$ be a class of real-valued functions and $l:\mathcal{F}\times\mathcal{Z}\times\mathcal{Y}\to\mathds{R}$ be a bounded loss function. We assume $Z\perp G$ and denote the rate of observed values as $m_{1}:=\mathds{E}\left[{\xi_{1}}\right]=rm_{1}^{1}+(1-r)m_{1}^{0}$.
Lemma 1 (Augmented loss).
Assume fixed $f\in\mathcal{F},z\in\mathcal{Z}$, $y\in\mathcal{Y}$ and $X\in\mathds{R}^{d}$ defined by $X_{1}=Z_{1}\xi_{1}$ and $X_{[2:d]}=z_{[2:d]}$. Define
$\tilde{l}(f,X,y):=\frac{1}{m_{1}}l(f,X,y)-\frac{1-m_{1}}{m_{1}}l(f,[0,X_{[2:d]}]^{T},y).$
If $Z\bot G$, we have that $\mathds{E}_{\xi_{1}}\left[\tilde{l}(f,X,y)\right]=l(f,z,y)$.
The fact that the augmented loss is unbiased with respect to under-reporting noise implies that a prediction model on observed data estimated with augmented loss, i.e. $\hat{f}=\arg\min_{f\in\mathcal{F}}\mathds{E}_{(X,Y)}\left[\tilde{l}(f,X,Y)\right],$ asymptotically recovers the Bayes optimal model on the true features $Z$. If $Y=\alpha+\beta^{T}Z$, $\mathcal{F}$ is the class of linear functions $f:\mathds{R}^{d}\to\mathds{R}$ and $l(f,z,y)=(f(z)-y)^{2}$ denotes squared error loss, the true parameters $\hat{\alpha}=\alpha$ and $\hat{\beta}=\beta$ are retrieved. Note that squared error loss is not bounded and estimating $\hat{f}$ requires the additional constraint $\tilde{l}(f,X,Y)\geq 0$. Lemma 1 operates in a group-agnostic setting with $Z\bot G$. Lemma A.3 provides a more general group-dependent version of the finding.
Optimal prediction imputation value
Assume we are in the linear case with $Y=\alpha+\beta^{T}Z$ and we have access to the true parameters $\alpha$ and $\beta$, e.g. obtained via augmented loss. What is the best possible prediction for an example of the form $x=[0,z_{2},\ldots,z_{d}]$? Since $x_{1}=0$ could mean $z_{1}=0$ or the entry is missing, it is intuitive that $\hat{y}=\alpha+\beta^{T}x$ does not minimize expected prediction error. Instead, we derive the optimal fixed prediction imputation value $x_{1}^{\prime\ast}$.
Lemma 2 (Optimal prediction imputation value).
Assume $Z\bot G$, $f(Z)=\alpha+\beta^{T}Z$ is the ground truth model and $X$ the random vector of observed features. We set
$X^{\prime}=\begin{cases}X\text{ if }X_{1}\neq 0,\\ [x_{1}^{\prime},X_{[2:d]}]\text{if }X_{1}=0,\end{cases}$
where $x_{1}^{\prime}$ is fixed. Then, $x_{1}^{\prime\ast}:=\arg\min_{x_{1}^{\prime}}\mathds{E}_{X}[(f(X^{\prime})-Y)^{2}]=\mathds{E}\left[{Z_{1}\mid X_{1}=0}\right]$ is the optimal prediction imputation value.
The Lemma shows that the loss-minimizing constant imputation value is the conditional mean $Z_{1}$ given the observed value is 0. This implies that, in alignment with earlier intuition, directly predicting with the observed $x=[0,z_{2},\ldots,z_{d}]$ is sub-optimal in the under-reporting setting. The optimal value $x_{1}^{\prime\ast}$ in the setting of Lemma 2 can be written as
$\mathds{E}\left[{Z_{1}\mid X_{1}=0}\right]=\frac{\frac{1}{m_{1}}\mathds{E}\left[{X_{1}}\right]-P(X_{1}\neq 0)\mathds{E}\left[{X_{1}\mid X_{1}\neq 0}\right]}{P(X_{1}=0)},$
which can be estimated directly from observed data if the under-reporting rate $1-m$ is known. If feature distributions vary across groups, group-dependent optimal prediction imputation values can be derived as described in Lemma A.4.
Under-reporting rate estimation
Both augmented loss estimation and optimal prediction imputation require access to the under-reporting rate $1-m$ which is typically unknown. In some cases, it may be possible to obtain supplementary data that can be used to estimate $m$. For example, in administrative data with under-reported health features for privately insured individuals, an external private insurance health claims dataset could be used to estimate the expected rate of true 0s. In most settings, estimation of under-reporting rates needs to rely directly on the observed data. Assume we have access to a dataset $V=\{(x,y)_{i=1}^{n}\}$. We split $V$ into a training portion $V_{\text{train}}$ and evaluation portion $V_{\text{eval}}$. Let $P$ denote the subset of examples from $V_{\text{eval}}$ for which $x_{1}\neq 0$. We draw on the literature on Positive and Unlabeled (PU) learning (Elkan and Noto, 2008) and estimate under-reporting rates as follows. First, we fit a model $h$ on $V_{\text{train}}$ to estimate $P(X_{1}\neq 0\mid X_{[2:d]},Y)$. Second, we evaluate $h$ on $P_{\text{eval}}$. The estimator for the share of observed values $m$ is given by $\hat{m}=\frac{1}{\mid P_{\text{eval}}\mid}\sum_{(x,y)\in P_{\text{eval}}}h(x_{[2:d],y}).$ The estimation procedure assumes that under-reporting occurs completely at random. Our experimental setting assumes under-reporting completely at random within groups and thus $m_{0}$ and $m_{1}$ can be estimated with the described procedure by restricting $V$ to examples from the respective group.
7. Experiments
7.1. Publicly available datasets
Data
Both COMPAS data (Angwin et al., 2016) and German credit data (Repository, 1994) are widely used across the algorithmic fairness literature. The American Community Survey (ACS) Income dataset is comprised of 2018 census data from California queried using the folktables package (Ding et al., 2021). The datasets vary in size, number of features, and prediction tasks as shown in Table D.1. We conduct all experiments with both gender and race as group columns if available. All models are group-blind and race information is never included as predictive feature. While the results for ACS Income and COMPAS are covered in the main text, we refer to Appendix F.1 for a discussion of the German credit data experiments.
Semi-synthetic outcomes
Since all of the prediction tasks are binary classification, we opt to generate semi-synthetic regression labels for our experiments. We first fit a logistic regression model to the entire dataset and extract the fitted probabilities. For the ACS Income data, the probabilities are rescaled to more closely reflect the \$50,000 income threshold. Next, we fit a linear regression model using the same features and the predicted probabilities as outputs. The fitted values from this linear model are chosen as the new “true” labels for our experiment. This outcome augmentation procedure allows us to generate artificial settings with a truly linear ground truth similar to the settings studied in Sections 4 and 5 while leaving realistic covariance structures intact. We further experiment with controlling the $R^{2}$ of the true linear model by adding additional noise to outcomes and report fairness implications in Appendix F.2.
Experiment stratification
Regression models are trained to predict semi-synthetic outcomes based on the features of the respective dataset. We then select the top $C$ share of the predictions as high risk and evaluate excess selection rates as given in Definition 1 to assess the fairness impact of under-reporting. Artificial under-reporting is added to one feature column at a time and we repeat the experiments for each outcome column, group column, and under-reporting rate. Under-reporting rates range from 0-90% in 10 percentage point increments, and we add under-reporting to only one group at a time (e.g., we set 10% of a feature in the male group to 0 while leaving the features of the female group unchanged). Only numeric features are considered for under-reporting since, in administrative data, binary features are often categorical dummies or thresholded versions of continuous count features. All models are trained with 80% of the datasets while withholding 20% for testing. In addition to the baseline models, we experiment with various solution approaches as described in Section 6. This includes our proposed methods of group-dependent augmented loss estimation and group-dependent optimal prediction imputation.
7.2. County-level birth data
Data
We present an analysis of a private administrative dataset we obtained from a county in the US. The dataset contains information on newborn children and their families including demographics, child protective services history, birth record data, and mental and behavioral health information for those who used publicly funded services. We set up a prediction task that attempts to mirror the analysis described in the Hello Baby model methodology report from Allegheny County (for Social Data Analytics at the Auckland University of Technology, 2020). The Hello Baby model was developed to predict which families are at greatest risk of having their child removed by Child Protective Services (CPS) during their first three years of life, and is used to prioritize families for opt-in, voluntary supportive services. Using our data, we train a similar model, and explore the effect of adding additional under-reporting to the behavioral and mental health data fields.
Experiment setup
We use the birth dataset with its original prediction outcomes to showcase a realistic example of the effect of under-reporting. As before, the data is separated into 80% for training and 20% for testing. We fit separate logistic regression models on three datasets. (1) Data as observed. (2) Data with behavioral health features set to 0 for privately insured individuals, i.e. mothers that are not insured through Medicaid. (3) Data without behavioral health features. For illustration, results are stratified by whether individuals are covered by Medicaid, and by whether individuals identify as African-American. Medicaid coverage and race are not used as features in any of the models.
8. Results
[…]
Figure 2. Group-wise excess selection rates using the ACS Income dataset. Each panel represents a feature that has been corrupted by under-reporting in independent runs of the experiment. Black curves show performance when omitting the entire feature column. Results are averaged over 50 runs on the test set. Shaded areas correspond to one standard deviation in each direction of the mean.
ACS Income data
Figure 2 summarizes the results for the experiments on the ACS Income data. We see that feature under-reporting in ‘education attainment’ and ‘hours worked per week’ consistently leads to under-selection of the group with under-reporting. This is true irrespective of whether feature under-reporting is injected into the female sub-group or the male sub-group, and we observe the same effect when under-reporting is added based on the individuals’ racial group. The figure additionally suggests that more under-reporting generally leads to increasing under-selection. Intuitively, it makes sense that both education attainment and hours worked per week contribute positively to predicted income which is confirmed by the parameter estimates (Figure D.3). Exploration of the covariance matrix of the unbiased features further reveals that all numeric columns in the dataset are positively correlated which together creates a setting reminiscent of the Case 2 scenario studied in Section 5. On a high-level, the scenario predicts that the group with more under-reporting is under-selected in the given setting which aligns with our observations. In addition to selection rate disparity, feature under-reporting in the data also leads to decreased model accuracy as displayed in Figure D.2, and the parameter estimates in Figure D.3 display an attenuation effect as predicted in Section 4.
COMPAS data
We focus on results for under-reporting in count features (Figure D.1) and point to Appendix F.3 for additional results. The feature ‘priors count’, i.e. the number of previous criminal offenses individuals have been convicted of, emerges as important feature with respect to under-reporting. Under-reporting in priors count leads to under-selection of the impacted group. This pattern repeats itself for any of the groups and both of the available prediction outcomes. The more feature under-reporting in a group, the larger the occurring outcome disparity. Similarly to the previous results, this suggests a setting of Case 2 as discussed in Section 5. As before, parameter estimates suggest an attenuation effect which is displayed in Figure 3. Under-reporting in priors count could be interpreted as an extreme case of crimes that do not result in arrest. Assuming that one demographic group is more likely to be convicted for committed crimes than the other group, the result implies that the already more frequently targeted group may additionally be flagged as high risk for recidivism at disproportionate rates. Racial disparities in arrest rates and police encounters are well-documented in the US (e.g. Alexander, 2010; Butcher et al., 2022; Fogliato et al., 2021; Pierson et al., 2020) which highlights the importance of this finding.
Standard missing data methods
Our experiments reveal that none of the standard missing data methods reliably mitigate the bias introduced through under-reporting and, instead, can have significant effects on excess selection rates themselves. Omission of the feature ‘h worked per week’ leads to over-selection of the female group and under-selection of the male group in the ACS Income data. This is because female individuals report to work on average less than male individuals (35.43h/week vs. 40.05h/week) while work hours contribute positively to income (Figure D.3). Omitting the feature blinds the model to these differences. Similar effects occur with the feature ‘education attainment’, and ‘priors count’ in the COMPAS data (Figures D.1 and 3).
[…]
Figure 3. Excess selection rates of group Other (i.e. not African-American) (left columns), parameter estimates (middle column), and test set $R^{2}$ (right columns) when under-reporting is injected into ‘priors count’ in group Other using the COMPAS dataset and synthetic two-year recidivism outcomes. In (a), the model is trained and evaluated using the under-reported feature. For (b), we first train a multiple imputation model and then train and evaluate the prediction model using probabilistic imputations. For (c), the model is trained on only rows without 0-entries in ‘priors count’ and evaluated on the under-reported data. In (d), we train with group-dependent augmented loss and use group-dependent optimal imputation values for prediction. Results are reported as averages over 30 runs. Shaded areas correspond to one standard deviation. The solid dots in the middle column correspond to true parameters. Note that in order to preserve readability, parameter estimates are only displayed for continuous features. Figure D.6 provides an overlay plot of the rightmost column for easy comparison.
For multiple imputation on the COMPAS data, we see that the excess selection rate flips signs and the group with under-reported ‘priors count’ is over-selected (Figures 3 and D.4). This is because the feature has a lot of true 0-entries that are wrongfully imputed as positive values. The cost incurred by these wrong imputations exceeds the benefit of imputation. In comparison to training on mismeasured features directly, the parameter estimation bias is considerable even for small amounts of under-reporting. For high under-reporting rates, the excess selection rate follows a similar pattern as the excess selection rate with feature omission since imputation is conducted using the features already present in the model adding little to no additional information. Since our models are well-specified, row omission recovers the true parameter estimates as displayed for the COMPAS data and feature ‘priors count’ in Figure 3. Despite access to the ground-truth, we observe that under-reporting bias introduced at prediction time increases the selection rate disparity. While the model without row omission is able to shift weight to correlated features as more and more entries for ‘priors count’ are under-reported, the row omission model cannot make use of the feature correlations ultimately leading to the increasing rather than decreasing disparities. With the same reasoning, the test set performance as measured by $R^{2}$ is decreased as displayed in the figure.
Figure 4. Selection rate fractions of different models. On the left, the results are displayed for the sub-population of Black individuals. On the right, the results are displayed for the sub-population that is insured through Medicaid. The selection rate of the whole population is considered to be 10% or lower which reflects a realistic range for predictive risk modeling.
Augmented loss and optimal prediction imputation
We contrast the performance of our method and standard approaches for handling missing features at the example of the ‘priors count’ feature in the COMPAS dataset. Under-reporting rates are estimated with the procedure described in Section 6. We refer to Appendix E for further details on the under-reporting rate estimation. The results in Figure 3 show that selection rate disparities decrease considerably when using group-dependent augmented loss and group-dependent optimal prediction imputation. In contrast multiple imputation, this fairness improvement comes at no visible cost in performance. In fact, the average test set $R^{2}$ of the corrected model is very similar to, and even slightly higher than, the test set $R^{2}$ of the model trained directly on under-reported data (see Figure D.6). Despite some variability, the average parameter estimates of the corrected model appear more stable across different amounts of under-reporting which suggests that the method successfully diminishes the bias introduced by under-reporting.
County-level birth data
For the birth data, Figure 4 suggest that under-reporting of all behavioral health data for the non-Medicaid population leads to over-selection of the Medicaid population. In the displayed overall selection rate range (¡10%), the Medicaid population is selected about 10% more often than in the “true” data setting. Note that in reality this difference could be even larger because some of the “true” data features were likely already under-reported. As shown in the Figure, some of the resulting disadvantage is still observable when evaluating performance in the group of Black individuals. This can be explained by the fact that the two group variables are positively correlated in the dataset ($\rho=0.44$). Removing the behavioral health columns altogether leads to a reversal of selection rate disparities in the Medicaid / non-Medicaid groups, and significantly increases selection of the Black sub-population. This gives further evidence that feature omission is not a successful remedy for the disparities caused by differential under-reporting.
9. Discussion
Differential feature under-reporting is a common phenomenon in administrative data. Data records are generally more complete for individuals who rely more consistently on public services (e.g. public health coverage). In many predictive risk assessment settings, the segment of the population with more complete observations overlaps with sub-populations that are more commonly selected as high-risk. Critics have argued that differential data availability is a key driver of the observed disparity in selection rates. When being classified as high risk subjects one to greater scrutiny of burden, this may disadvantage those with more complete data (Eubanks, 2018). Overall, the results of our study lend further credence to the concern by demonstrating how feature under-reporting generally leads to under-selection of a group that is already less frequently identified as high-risk. While, as we demonstrate, groups with greater data availability can theoretically be under-selected, the feature dependence structure under which this occurs appears to be uncommon in practice.
We illustrate the increased selection rate that individuals who rely on public healthcare coverage may experience at a real world example. Following the idea of Allegheny County’s Hello Baby program (for Social Data Analytics at the Auckland University of Technology, 2020), we build a model that predicts the risk that a newborn child will be removed from their family by Child Protective Services (CPS) within three years based on county-level data. The dataset contains behavioral and mental health information on the parents which can be assumed to be more complete for families that rely on public insurance. We note that, for privately insured individuals, some of this information may still be observed, e.g. because the individual was publicly insured previously, or individual information has been collected explicitly, but a lot of the information can be assumed missing. Our experiments suggest that further under-reporting in behavioral health related information for the privately insured sub-population leads to an increase in high-risk predictions for the publicly insured group. We hypothesize that this effect would be even larger if the dataset was not already missing large portions of the feature observations for the privately insured group. This finding implies an unfair targeting of the publicly insured sub-population as high-risk. Since African-American mothers in the data are publicly insured more frequently than mothers from other racial groups, these results also suggest that the newborns of African-American mothers are predicted to be at high-risk at unfairly inflated rates. Of course, depending on intervention type, high risk may lead to an advantage or disadvantage for the families. In the Hello Baby setting, it is tied to eligibility for voluntary supportive services provided by county-funded service providers.
Our work proposes a technical remedy for the impact of under-reporting as a driver of disparities in selection rates. While standard missing data methods did not lead to more equitable outcomes in our experiments, these new methods reduced disparities considerably with little to no decrease in model accuracy. The applicability and performance in real-world administrative data settings like the Hello baby program remains an interesting and important avenue for future work.
Research ethics and social impact
Ethical considerations statement
The authors did not face ethical concerns that had to be mitigated while conducting this study. Experiments in this paper are based on the commonly used COMPAS and German credit datasets, 2018 US Census data from the American Community Survey, and a private county level data set. Result are aggregated over broad population groups and no identifiable information can be retrieved.
Researcher positionality statement
The authors recognize that their societal advantages give them certain benefits not shared by all individuals undergoing public sector risk assessment. Thus, great care was taken in reflecting on the question: Does this work benefit us or the community at large? Since inflated selection rates can lead to tangible advantages (e.g. qualification for further publicly funded services) and significant disadvantages (e.g. unfavorable bail decisions in criminal risk assessment) depending on the application area, we believe that addressing problems of feature under-reporting ultimately benefits the community at large.
Adverse impact statement
The authors believe that drawing attention to the problem of differential feature under-reporting has the potential to positively impact public sector risk assessment instruments for all individuals subjected to these systems. However, our work provides only a first step towards finding appropriate solutions to this problem. We propose a potential mitigation method and evaluate the method in a semi-synthetic setting, but assessment of potential adverse effects of the method in real-world applications are beyond the scope of this paper. We clearly state this at the end of the discussion and call for future work in this direction.
Appendix A Additional theorems
Proposition A.1 ().
In the $d$-dimensional case setting described in Section 4, the parameter estimates from Equation 1 take the form
(2) $\begin{split}\hat{\beta}_{1}&=\beta_{1}\frac{1}{1-R^{2}}\sqrt{\frac{\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{X_{1}}\right]}}\left(\rho(X_{1},Z_{1})-\sum_{i=2}^{d}\rho(X_{1},Z_{i})\rho(Z_{1},Z_{i})\right),\\ \hat{\beta}_{k}&=\beta_{1}\sqrt{\frac{\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{k}}\right]}}\left(\rho(Z_{k},Z_{1})-\frac{1}{1-R^{2}}\rho(X_{1},Z_{k})\left(\rho(X_{1},Z_{1})-\sum_{i=2}^{d}\rho(X_{1},Z_{i})\rho(Z_{1},Z_{i})\right)\right)+\beta_{k}\\ \end{split}$
for $k\in[2:d]$. Here, $R^{2}=\sum_{i=2}^{d}\rho(X_{1},Z_{i})^{2}\in[0,1)$.
Here, $R^{2}$ is the squared coefficient of multiple correlation between $Z_{1}\xi_{1}$ and $Z_{[2:d]}=[Z_{2},\ldots,Z_{d}]$ which can be interpreted as the fraction of variance in $Z_{1}\xi_{1}$ that can be explained by the independent variables $Z_{[2:d]}$. If all features are observed, the factor $\rho(X_{1},Z_{1})-\sum_{i=2}^{d}\rho(X_{1},Z_{i})\rho(Z_{1},Z_{i})$ collapses to $(1-R^{2})$, and the estimates are unbiased. With under-reporting the bias introduced into the parameter estimates depends on the strength of correlations between features, as well as how this correlation changes with the mismeasurement of $Z_{1}$. The bias in Equation 2 can be conceptualized as a generalization of omitted variable bias (Angrist and Pischke, 2008) which is further explored in Appendix B.
Proposition A.2 ().
Define the threshold turning point $T$ as
[…]
Then, for a high threshold $\tilde{y}$ with $\tilde{y}>T$, the group with more under-reporting is
  • Case 1: Over-selected if $\mathds{V}\left[{\hat{\alpha}+\hat{\beta}_{[2:d]}Z_{[2:d]}}\right]>\mathds{V}\left[{\hat{\alpha}+\hat{\beta}Z}\right]$, or
  • Case 2: Under-selected if $\mathds{V}\left[{\hat{\alpha}+\hat{\beta}_{[2:d]}Z_{[2:d]}}\right]<\mathds{V}\left[{\hat{\alpha}+\hat{\beta}Z}\right]$.
For low thresholds $\tilde{y}<T$, the cases are reversed.
In practical applications, thresholds are usually set such that only a small portion of predictions exceeds the threshold. For example, we can only decide to screen a small portion of calls in the child welfare setting. In particular, realistic thresholds are generally well above the average $\hat{Y}$. On a high level, the turning point $T$ in Proposition A.2 represent an adjusted mean predicted value where the influence of the feature with under-reporting is weighed depending on a ratio determined by prediction variances with and without the feature.
Lemma A.3 (Group-dependent augmented loss).
Assume fixed $f\in\mathcal{F},z\in\mathcal{Z},y\in\mathcal{Y},g\in\{0,1\}$ and $X\in\mathds{R}^{d}$ defined by $X_{1}=Z_{1}\xi_{1}^{g}$ and $X_{[2:d]}=z_{[2:d]}$. Define
$\tilde{l}(f,X,y,g)$ $:=\frac{1}{m_{1}^{g}}l(f,X,y)-\frac{1-m_{1}^{g}}{m_{1}^{g}}l(f,[0,X_{[2:d]}]^{T},y)$
Then, we have that $\mathds{E}_{\xi_{1}^{g}}\left[\tilde{l}(f,X,y,g)\right]=l(f,z,y)$.
Lemma A.4 (Group-dependent optimal prediction imputation values).
Assume $f(Z)=\alpha+\beta^{T}Z$ is the ground truth model, $X$ a random vector of observed features, and $G$ the group membership. We set
$X^{\prime}=\begin{cases}X\text{ if }X_{1}\neq 0,\\ [x_{1}^{\prime 0},X_{[2:d]}]\text{ if }X_{1}=0\text{ and }G=0,\\ [x_{1}^{\prime 1},X_{[2:d]}]\text{ if }X_{1}=0\text{ and }G=1,\\ \end{cases}$
where $x_{1}^{\prime 0},x_{1}^{\prime 1}$ are group-dependent fixed imputation values. Then,
$\arg\min_{x_{1}^{\prime g}}\mathds{E}_{X}\left[(f(X^{\prime})-Y)^{2}\right]=\mathds{E}\left[{Z_{1}\mid X_{1}=0,G=g}\right]$
are the optimal group-dependent prediction imputation values for $g\in\{0,1\}$.
Similar to before, the optimal imputation values can be written as
$\mathds{E}\left[{Z_{1}\mid X_{1}=0,G=g}\right]=\frac{\frac{1}{m_{1}^{g}}\mathds{E}\left[{X_{1},G=g}\right]-P(X_{1}\neq 0\mid G=g)\mathds{E}\left[{X_{1}\mid X_{1}\neq 0,G=g}\right]}{P(X_{1}=0\mid G=g)},$
which can be estimated directly from observed data.
Appendix B Connection between Proposition A.1 and omitted variable bias
Econometrics literature uses the term omitted variable bias to refer to the model estimation bias that is introduced when omitting an independent variable that influences both other independent variables and the dependent outcome (Angrist and Pischke, 2008). In the setting of Proposition A.1, omitting the first feature entirely corresponds to a setting in which all feature entries are under-reported, i.e. default to 0. The $k$-th parameter estimate in this case can be written as
$\hat{\beta}_{k}=\beta_{1}\frac{\text{Cov}\left[{Z_{k}},{Z_{1}}\right]}{\mathds{V}\left[{Z_{k}}\right]}+\beta_{k}$
which is known as omitted variable bias formula (Angrist and Pischke, 2008). Here, $\gamma_{Z_{1},Z_{k}}=\text{Cov}\left[{Z_{1}},{Z_{k}}\right]/\mathds{V}\left[{Z_{k}}\right]$ corresponds to the population regression coefficient of a linear regression of $Z_{1}$ on $Z_{k}$ which can be written as
$Z_{1}=\alpha_{Z_{1},Z_{k}}+\gamma_{Z_{1},Z_{k}}Z_{k},$
where $\alpha_{Z_{1},Z_{k}}$ is an intercept. Omitting $Z_{1}$ from the regression induces a confounding relationship where the effects of $Z_{1}$ on $Z_{k}$ become intertwined. Instead of isolating the effect of $Z_{k}$ on $Y$, $\hat{\beta}_{k}$ also includes a partial effect of $Z_{1}$ on $Y$. This effect is scaled by $\gamma_{Z_{1},Z_{k}}$ to account for the linear relationship between $Z_{1}$ and $Z_{k}$.
In the setting of this paper, we are interested in cases in which some but not necessarily all of the feature entries are missing. Maintaining the same notation as before, $\hat{\beta}_{1}$ from Equation 2 in this general case can be written as
$\hat{\beta_{1}}$ $=\beta_{1}\frac{1}{1-R^{2}}\left(\frac{\text{Cov}\left[{X_{1}},{Z_{1}}\right]}{\mathds{V}\left[{X_{1}}\right]}-\sum_{i=2}^{d}\frac{\text{Cov}\left[{X_{1}},{Z_{i}}\right]\text{Cov}\left[{Z_{1}},{Z_{i}}\right]}{\mathds{V}\left[{X_{1}}\right]\mathds{V}\left[{Z_{i}}\right]}\right)$
$=\beta_{1}\left(\frac{\gamma_{Z_{1},X_{1}}-\sum_{i=2}^{d}\gamma_{Z_{i},X_{1}}\gamma_{Z_{1},Z_{i}}}{1-R^{2}}\right).$
Here, the numerator of the biasing factor reflects how much information about $Z_{1}$ remains encoded in $X_{1}$ without drawing on associations through the other features $Z_{2},\ldots,Z_{d}$ (i.e, arrows of the form $X_{1}\to Z_{i}\to Z_{1}$). The denominator measures how much of the variance in $X_{1}$ is explained by $Z_{2},\ldots,Z_{d}$. For $k\in[2:d]$, we receive
$\hat{\beta}_{k}$ $=\beta_{k}+\beta_{1}\frac{\text{Cov}\left[{Z_{k}},{Z_{1}}\right]}{\mathds{V}\left[{Z_{k}}\right]}-\beta_{1}\frac{1}{1-R^{2}}\frac{\text{Cov}\left[{X_{1}},{Z_{k}}\right]}{\mathds{V}\left[{Z_{k}}\right]}\left(\frac{\text{Cov}\left[{X_{1}},{Z_{1}}\right]}{\mathds{V}\left[{X_{1}}\right]}-\sum_{i=2}^{d}\frac{\text{Cov}\left[{X_{1}},{Z_{i}}\right]\text{Cov}\left[{Z_{1}},{Z_{i}}\right]}{\mathds{V}\left[{Z_{i}}\right]\mathds{V}\left[{X_{1}}\right]}\right)$
$=\beta_{k}+\beta_{1}\gamma_{Z_{1},Z_{k}}-\beta_{1}\frac{1}{1-R^{2}}\gamma_{X_{1},Z_{k}}\left(\gamma_{Z_{1},X_{1}}-\sum_{i=2}^{d}\gamma_{Z_{i},X_{1}}\gamma_{Z_{1},Z_{i}}\right)$
$=\beta_{k}+\underbrace{\beta_{1}\gamma_{Z_{1},Z_{k}}}_{\text{omitted variable bias}}-\underbrace{\hat{\beta}_{1}\gamma_{X_{1},Z_{k}}.}_{\text{Correction since partially observed}}$
Instead of just encoding the effect of $Z_{k}$ on $Y$ and partial effect of $Z_{1}$ on $Y$ like before, the estimate $\hat{\beta}_{k}$ now also corrects for the fact that $Z_{1}$ is partially observed. The magnitude of the correction depends on the parameter estimate for the partially observed variable as well as the linear relationship between $X_{1}$ and $Z_{k}$.
Appendix C Proofs
In this section, we provide the full proofs for the results in the main text.
Lemma 1
We have $\xi\perp Z$ and $\mathds{E}\left[{\xi^{2}}\right]=\mathds{E}\left[{\xi}\right]$. Since $\mathds{E}\left[{\xi}\right]\in[0,1]$, we have
$\mid\hat{\beta}\mid=\mid\frac{\text{Cov}\left[{X},{Z}\right]}{\text{Cov}\left[{X},{X}\right]}\beta\mid=\mid\frac{\mathds{E}\left[{\xi Z^{2}}\right]-\mathds{E}\left[{\xi Z}\right]\mathds{E}\left[{Z}\right]}{\mathds{E}\left[{\xi^{2}Z^{2}}\right]-\mathds{E}\left[{\xi Z}\right]^{2}}\beta\mid=\frac{\mathds{E}\left[{\xi}\right]\mathds{V}\left[{Z}\right]}{\mathds{E}\left[{\xi}\right](\mathds{E}\left[{Z^{2}}\right]-\mathds{E}\left[{\xi}\right]\mathds{E}\left[{Z}\right]^{2})}\mid\beta\mid\leq\mid\beta\mid.$
Proposition A.1
In order to derive the equations for $\hat{\beta}_{i}$ for $i\in[1:d]$, we start by inverting the covariance matrix
$\Sigma_{X}=\begin{pmatrix}\mathds{V}\left[{Z_{1}\xi_{1}}\right]&\text{Cov}\left[{Z_{1}\xi_{1}},{Z_{2}}\right]&\text{Cov}\left[{Z_{1}\xi_{1}},{Z_{3}}\right]&\cdots&\text{Cov}\left[{Z_{1}\xi_{1}},{Z_{d}}\right]\\ \text{Cov}\left[{Z_{2}},{Z_{1}\xi_{1}}\right]&\mathds{V}\left[{Z_{2}}\right]&\text{Cov}\left[{Z_{2}},{Z_{3}}\right]&\cdots&\text{Cov}\left[{Z_{2}},{Z_{d}}\right]\\ \text{Cov}\left[{Z_{3}},{Z_{1}\xi_{1}}\right]&\text{Cov}\left[{Z_{3}},{Z_{2}}\right]&\mathds{V}\left[{Z_{3}}\right]&\cdots&\text{Cov}\left[{Z_{3}},{Z_{d}}\right]\\ \vdots&&&\ddots&\vdots\\ \text{Cov}\left[{Z_{d}},{Z_{1}\xi_{1}}\right]&\text{Cov}\left[{Z_{d}},{Z_{2}}\right]&\cdots&&\mathds{V}\left[{Z_{d}}\right]\end{pmatrix}.$
For this, we separate the matrix into the blocks $A=(\Sigma_{X})_{11}$, $B=((\Sigma_{X})_{1j})_{j\in[2:d]}$, $C=((\Sigma_{X})_{i1})_{i\in[2:d]}$, and $D=((\Sigma_{X})_{ij})_{i,j\in[2:d]}$. Note that $A\in\mathds{R}^{1\times 1}$, $B=C^{T}\in\mathds{R}^{1\times(d-1)}$, and $D\in\mathds{R}^{(d-1)\times(d-1)}$. Using matrix inversion theorem, the inverse of $\Sigma_{X}$ can be written as
(3) $\begin{split}(\Sigma_{X})^{-1}=\begin{pmatrix}A^{-1}+A^{-1}B(D-CA^{-1}B)^{-1}CA^{-1}&-A^{-1}B(D-CA^{-1}B)^{-1}\\ -(D-CA^{-1}B)^{-1}CA^{-1}&(D-CA^{-1}B)^{-1}\end{pmatrix},\end{split}$
where $D-CA^{-1}B=D-CA^{-1}C^{T}$ is the Schur complement of $A$ in $\Sigma_{X}$. Recall that, by assumption, $\text{Cov}\left[{Z_{i}},{Z_{j}}\right]=0$ for $i,j>1$ with $i\neq j$ which means that $D$ is a diagonal matrix. We also note that $\text{rank}(CC^{T})=1$. Denoting $g=\text{trace}(-A^{-1}CC^{T}D^{-1})$, the inverse of the Schur complement can be written as
$(D-CA^{-1}C^{T})^{-1}=(D-A^{-1}CC^{T})^{-1}=D^{-1}+\frac{1}{1+g}D^{-1}A^{-1}CC^{T}D^{-1}.$
Here, $D^{-1}$ is a diagonal matrix with values $1/\mathds{V}\left[{Z_{i}}\right]$ for $i\in[2:d]$ on the diagonal, $A^{-1}=1/\mathds{V}\left[{Z_{1}\xi_{1}}\right]$, and the diagonal of $CC^{T}$ is $\text{Cov}\left[{Z_{1}\xi_{1}},{Z_{i}}\right]^{2}$ for $i\in[2:d]$. It follows that
$g=-\sum_{i=2}^{d}\rho(Z_{1}\xi_{1},Z_{i})^{2}$
which corresponds to the negative of the $R^{2}$ between $Z_{1}\xi$ and $Z_{2},\ldots,Z_{d}$. We hence write $R^{2}$ for $-g$ in the following.
Now we can calculate the top left block of the inverse matrix in Equation 3 as
$A^{-1}+A^{-1}B(D-CA^{-1}B)^{-1}CA^{-1}=\frac{1}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\frac{1}{1-R^{2}}.$
The top right bock corresponds to the row vector
$-AB(D-CA^{-1}B)^{-1}=\left(\left(-\frac{\text{Cov}\left[{Z_{1}\xi_{1}},{Z_{i}}\right]}{\mathds{V}\left[{Z_{i}}\right]\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\frac{1}{1-R^{2}}\right)_{i}\right)_{i\in[2:d]},$
while the bottom left block is the same transposed. Lastly, the bottom left block of Equation 3 can be computed as
$(D-CA^{-1}B)^{-1}=\left(\text{diag}(1/\mathds{V}\left[{Z_{i}}\right])+\frac{1}{1-R^{2}}\left(\frac{\text{Cov}\left[{Z_{1}\xi_{1}},{Z_{i}}\right]\text{Cov}\left[{Z_{1}\xi_{1}},{Z_{j}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]\mathds{V}\left[{Z_{i}}\right]\mathds{V}\left[{Z_{j}}\right]}\right)_{ij}\right)_{i,j\in[2:d]}.$
Inserting these values into Equation 1 yields the desired parameter estimates
$\hat{\beta}_{1}$ $=\beta_{1}\frac{1}{1-R^{2}}\sqrt{\frac{\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{X_{1}}\right]}}\left(\rho(X_{1},Z_{1})-\sum_{i=2}^{d}\rho(X_{1},Z_{i})\rho(Z_{1},Z_{i})\right),$
and
$\hat{\beta}_{k}$ $=\beta_{1}\sqrt{\frac{\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{k}}\right]}}\left(\rho(Z_{k},Z_{1})-\frac{1}{1-R^{2}}\rho(X_{1},Z_{k})\left(\rho(X_{1},Z_{1})-\sum_{i=2}^{d}\rho(X_{1},Z_{i})\rho(Z_{1},Z_{i})\right)\right)+\beta_{k}.$
Proposition 2
The estimates from Proposition A.1 simplify to
$\hat{\beta}_{1}=\frac{1}{1-R^{2}}\left(\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}-\frac{R^{2}}{\mathds{E}\left[{\xi_{1}}\right]}\right)\beta_{1}$
and
$\hat{\beta}_{k}=\beta_{1}\frac{1}{1-R^{2}}\left(\frac{\text{Cov}\left[{Z_{1}},{Z_{k}}\right](\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2})}{\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\right)+\beta_{k},$
for $k\in[2:d]$.
For the first claim, recall that $\mathds{V}\left[{\xi}\right]=\mathds{E}\left[{\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}$ and $\mathds{V}\left[{Z_{1}\xi_{1}}\right]=\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2}+\mathds{V}\left[{\xi_{1}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]$. Note that $Z\perp\xi$ allows us to rewrite
$R^{2}$ $=\sum_{i=2}^{d}\rho(Z_{1}\xi_{1},Z_{i})^{2}$
$=\sum_{i=2}^{d}\frac{\mathds{E}\left[{\xi_{1}}\right]^{2}\text{Cov}\left[{Z_{1}},{Z_{i}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]\mathds{V}\left[{Z_{i}}\right]\mathds{V}\left[{Z_{1}}\right]}$
$=\frac{\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}$
and thus
$\frac{1}{1-R^{2}}\left(\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}-\frac{R^{2}}{\mathds{E}\left[{\xi_{1}}\right]}\right)$
$=$ $\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]-R^{2}\mathds{V}\left[{Z_{1}\xi_{1}}\right]}{(1-R^{2})\mathds{V}\left[{Z_{1}\xi_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]}$
$=$ $\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]-\frac{\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}\mathds{V}\left[{Z_{1}\xi_{1}}\right]}{(1-\frac{\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2})\mathds{V}\left[{Z_{1}\xi_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]}$
$=$ $\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}}{(\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2}+(\mathds{E}\left[{\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2})\mathds{E}\left[{Z_{1}^{2}}\right])\mathds{E}\left[{\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}\mathds{E}\left[{\xi_{1}}\right]}$
$=$ $\frac{\mathds{V}\left[{Z_{1}}\right](1-\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2})}{\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right](1-\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2})+(1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]},$
which is positive as long as $Z_{1}$ is not a linear combination of other features which was explicitly excluded from consideration. The claim follows.
For the second claim, we show that
$\frac{\mathds{V}\left[{Z_{1}}\right](1-\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2})}{\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right](1-\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2})+(1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]}<1$
$\Leftrightarrow$ $(1-\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2})\mathds{V}\left[{Z_{1}}\right](1-\mathds{E}\left[{\xi_{1}}\right])<(1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]$
$\Leftrightarrow$ $(1-\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2})(\mathds{E}\left[{Z_{1}^{2}}\right]-\mathds{E}\left[{Z_{1}}\right]^{2})<\mathds{E}\left[{Z_{1}^{2}}\right]$
$\Leftrightarrow$ $(1-\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2})\left(1-\frac{\mathds{E}\left[{Z_{1}}\right]^{2}}{\mathds{E}\left[{Z_{1}}\right]^{2}}\right)<1.$
Since $Z_{1}$ is not a linear combination of other features, we know that $1-\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}\in(0,1]$ and this inequatily is always true. The claim follows with the first part of the proposition.
For the third claim, recall that $\mathds{E}\left[{\xi_{1}}\right]=rm^{1}_{1}+(1-r)m^{0}_{1}$. Assume we have two sets of parameters $(m_{1}^{0},m_{1}^{1})$ and $(m_{1}^{0^{\prime}},m_{1}^{1^{\prime}})$. If $m_{1}^{0}<m_{1}^{0^{\prime}}$ and $m_{1}^{1}\leq m_{1}^{1^{\prime}}$, we have
$\mathds{E}\left[{\xi}\right]=rm^{1}_{1}+(1-r)m^{0}_{1}<rm^{1^{\prime}}_{1}+(1-r)m^{0^{\prime}}_{1}=\mathds{E}\left[{\xi^{\prime}}\right].$
The same holds true if $m_{1}^{0}\leq m_{1}^{0^{\prime}}$ and $m_{1}^{1}<m_{1}^{1^{\prime}}$ which shows that the expected share of observed features $\mathds{E}\left[{\xi_{1}}\right]$ is decreasing if and only if we are increasing under-reporting in either (or both) of the groups while leaving everything else fixed. Thus, instead of changes in $m_{1}^{g}$, we argue directly about changes in $\mathds{E}\left[{\xi_{1}}\right]$ in the following.
Denote $S^{2}=\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}$ and consider the function
$f:(0,1]$ $\to\mathds{R}$
$\mathds{E}\left[{\xi_{1}}\right]=x$ $\mapsto\frac{\mathds{V}\left[{Z_{1}}\right](1-S^{2})}{\mathds{V}\left[{Z_{1}}\right]x(1-S^{2})+(1-x)\mathds{E}\left[{Z_{1}^{2}}\right]}.$
We show that $f$ is monotonically increasing from which the claim follows directly. It holds that
$\frac{d}{dx}f(x)=\frac{-\mathds{V}\left[{Z_{1}}\right](1-S^{2})(\mathds{V}\left[{Z_{1}}\right](1-S^{2})-\mathds{E}\left[{Z_{1}^{2}}\right])}{\left(\mathds{V}\left[{Z_{1}}\right]x(1-S^{2})+(1-x)\mathds{E}\left[{Z_{1}^{2}}\right]\right)^{2}}.$
Since $1-S^{2}\in[0,1)$ and $\mathds{V}\left[{Z_{1}}\right]>0$, the numerator is positive iff
$-\mathds{V}\left[{Z_{1}}\right](1-S^{2})(\mathds{V}\left[{Z_{1}}\right](1-S^{2})-\mathds{E}\left[{Z_{1}^{2}}\right])>0$
$\Leftrightarrow$ $\mathds{V}\left[{Z_{1}}\right](1-S^{2})<\mathds{E}\left[{Z_{1}^{2}}\right]$
$\Leftrightarrow$ $\left(1-\frac{\mathds{E}\left[{Z_{1}}\right]^{2}}{\mathds{E}\left[{Z_{1}^{2}}\right]}\right)(1-S^{2})<1,$
which is a true statement. We conclude that $f$ is monotonically increasing in $x$ and the claim follows.
Proposition 3
Given a $k\in[2:d]$, we know that
$\hat{\beta}_{k}=\beta_{1}\frac{1}{1-R^{2}}\left(\frac{\text{Cov}\left[{Z_{1}},{Z_{k}}\right](\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2})}{\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\right)+\beta_{k}.$
The first claim is obvious from this expression.
For the second claim, we follow similar steps as for the third claim in the proof of Proposition 2. We know from the previous proof that
$R^{2}=\frac{\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}$
and thus, denoting $S^{2}=\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}$,
$\frac{1}{1-R^{2}}\left(\frac{\text{Cov}\left[{Z_{1}},{Z_{k}}\right](\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{V}\left[{Z_{2}}\right]\mathds{E}\left[{\xi}\right]^{2})}{\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\right)$
$=$ $\frac{\text{Cov}\left[{Z_{1}},{Z_{k}}\right]\mathds{V}\left[{\xi_{1}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]}{(1-\frac{\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}\xi_{1}}\right]}$
$=$ $\frac{\text{Cov}\left[{Z_{1}},{Z_{k}}\right]\mathds{V}\left[{\xi_{1}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]}{\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]S^{2}\mathds{V}\left[{Z_{k}}\right]}$
$=$ $\frac{\text{Cov}\left[{Z_{1}},{Z_{k}}\right](\mathds{E}\left[{\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2})\mathds{E}\left[{Z_{1}^{2}}\right]}{\mathds{V}\left[{Z_{k}}\right](\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2}+(\mathds{E}\left[{\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2})\mathds{E}\left[{Z_{1}^{2}}\right])-\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]S^{2}\mathds{V}\left[{Z_{k}}\right]}$
$=$ $\frac{\text{Cov}\left[{Z_{1}},{Z_{k}}\right](1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]}{(1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]+\mathds{V}\left[{Z_{k}}\right](1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]},$
since $\mathds{V}\left[{Z_{1}\xi_{1}}\right]=\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2}+\mathds{V}\left[{\xi_{1}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]$ and $\mathds{V}\left[{\xi_{1}}\right]=\mathds{E}\left[{\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}$.
Now, consider the function
$g:(0,1]$ $\to\mathds{R}$
$\mathds{E}\left[{\xi_{1}}\right]=x$ $\mapsto\frac{\text{Cov}\left[{Z_{1}},{Z_{k}}\right](1-x)\mathds{E}\left[{Z_{1}^{2}}\right]}{(1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]x+\mathds{V}\left[{Z_{k}}\right](1-x)\mathds{E}\left[{Z_{1}^{2}}\right]}.$
We compute
$\frac{d}{dx}g(x)=$ $\frac{-\text{Cov}\left[{Z_{1}},{Z_{k}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]((1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]x+\mathds{V}\left[{Z_{k}}\right](1-x)\mathds{E}\left[{Z_{1}^{2}}\right])}{\left((1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]x+\mathds{V}\left[{Z_{k}}\right](1-x)\mathds{E}\left[{Z_{1}^{2}}\right]\right)^{2}}$
$-\frac{(1-x)\text{Cov}\left[{Z_{1}},{Z_{k}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]((1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]-\mathds{V}\left[{Z_{k}}\right]\mathds{E}\left[{Z_{1}^{2}}\right])}{\left((1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]x+\mathds{V}\left[{Z_{k}}\right](1-x)\mathds{E}\left[{Z_{1}^{2}}\right]\right)^{2}}.$
Further,
$\frac{d}{dx}g(x)>0$
$\Leftrightarrow$ $-\text{Cov}\left[{Z_{1}},{Z_{k}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]((1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]x+\mathds{V}\left[{Z_{k}}\right](1-x)\mathds{E}\left[{Z_{1}^{2}}\right])$
$>(1-x)\text{Cov}\left[{Z_{1}},{Z_{k}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]((1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]-\mathds{V}\left[{Z_{k}}\right]\mathds{E}\left[{Z_{1}^{2}}\right])$
$\Leftrightarrow$ $-\text{Cov}\left[{Z_{1}},{Z_{k}}\right](1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]x>(1-x)\text{Cov}\left[{Z_{1}},{Z_{k}}\right](1-S^{2})\mathds{V}\left[{Z_{k}}\right]\mathds{V}\left[{Z_{1}}\right]$
$\Leftrightarrow$ $0>\text{Cov}\left[{Z_{1}},{Z_{k}}\right].$
This shows that factor determining the influence of $\beta_{1}$ on $\hat{\beta}_{k}$ is increasing with decreasing under-reporting if $\text{Cov}\left[{Z_{1}},{Z_{k}}\right]<0$ and decreasing with decreasing under-reporting otherwise. The claim follows.
Proposition A.2
Predictions are obtained from the model $\hat{Y}=\hat{\alpha}+\hat{\beta}^{T}X$. Since $Z\sim\mathcal{N}(\mu,\Sigma)$ is jointly Gaussian, we know that
$\hat{\beta}^{T}Z\sim\mathcal{N}\left(\hat{\beta}^{T}\mu,\hat{\beta}^{T}\Sigma\hat{\beta}\right)=\mathcal{N}\left(\sum_{i=1}^{d}\hat{\beta}_{i}\mu_{i},\sum_{i=1}^{d}\hat{\beta}_{i}^{2}\sigma_{i}^{2}+\sum_{i=1}^{d}\sum_{j=i+1}^{d}2\hat{\beta}_{i}\hat{\beta}_{j}\text{Cov}\left[{Z_{i}},{Z_{j}}\right]\right)$
and
$\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}\sim\mathcal{N}\left(\hat{\beta}_{[2:d]}^{T}\mu_{[2:d]},\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}\right)=\mathcal{N}\left(\sum_{i=2}^{d}\hat{\beta}_{i}\mu_{i},\sum_{i=2}^{d}\hat{\beta}_{i}^{2}\sigma_{i}^{2}+\sum_{i=2}^{d}\sum_{j=i+1}^{d}2\hat{\beta}_{i}\hat{\beta}_{j}\text{Cov}\left[{Z_{i}},{Z_{j}}\right]\right)$
where $\sigma_{i}^{2}=\mathds{V}\left[{Z_{i}}\right]$ for $i\in[1:d]$.
The cdf of predictions $\hat{Y}$ in group $g\in\{0,1\}$ can be written as
$F_{\hat{Y}\mid G=g}(x)=$ $P\left(\hat{\beta}_{1}Z_{1}\xi_{1}^{g}+\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}\leq x-\hat{\alpha}\right)$
$=$ $(1-m_{1}^{g})P\left(\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}\leq x-\hat{\alpha}\right)+m_{1}^{g}P\left(\hat{\beta}^{T}Z\leq x-\hat{\alpha}\right).$
Let $C\in[0,1]$ and denote $\tilde{y}=F_{\hat{Y}}^{-1}(1-C)$. Without loss of generality, assume that $m_{1}^{0}<m_{1}^{1}$. If $m_{1}^{0}=m_{1}^{1}$ the selection rate disparity is 0, if $m_{1}^{0}>m_{1}^{1}$ the following calculation can easily be adjusted. The inequality $m_{1}^{0}<m_{1}^{1}$ means that group 0 has the same or more expected under-reporting in feature $Z_{1}$ than group 1. Group 0 is over-selected at threshold $C$ according to Definition 2 if and only if
$1-F_{\hat{Y}\mid G=1}(\tilde{y})<1-F_{\hat{Y}\mid G=0}(\tilde{y})$
$\Leftrightarrow$ $(1-m_{1}^{1})P\left(\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}\leq\tilde{y}-\hat{\alpha}\right)+m_{1}^{1}P\left(\hat{\beta}^{T}Z\leq\tilde{y}-\hat{\alpha}\right)>(1-m_{1}^{0})P\left(\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}\leq\tilde{y}-\hat{\alpha}\right)+m_{1}^{0}P\left(\hat{\beta}^{T}Z\leq\tilde{y}-\hat{\alpha}\right)$
$\Leftrightarrow$ $(m_{1}^{1}-m_{1}^{0})P\left(\hat{\beta}^{T}Z\leq\tilde{y}-\hat{\alpha}\right)>(m_{1}^{1}-m^{0}_{1})P\left(\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}\leq\tilde{y}-\hat{\alpha}\right)$
$\Leftrightarrow$ $P\left(\hat{\beta}^{T}Z\leq\tilde{y}-\hat{\alpha}\right)>P\left(\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}\leq\tilde{y}-\hat{\alpha}\right).$
Expanding on this in the jointly Gaussian case, we can see that
$P\left(\hat{\beta}^{T}Z\leq\tilde{y}-\hat{\alpha}\right)>P\left(\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}\leq\tilde{y}-\hat{\alpha}\right)$
$\Leftrightarrow$ $\Phi\left(\frac{\tilde{y}-\hat{\alpha}-\hat{\beta}^{T}\mu}{\sqrt{\hat{\beta}^{T}\Sigma\hat{\beta}}}\right)>\Phi\left(\frac{\tilde{y}-\hat{\alpha}-\hat{\beta}_{[2:d]}^{T}\mu_{[2:d]}}{\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}}\right)$
$\Leftrightarrow$ $\left(\tilde{y}-\hat{\alpha}-\hat{\beta}^{T}\mu\right)\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}>\left(\tilde{y}-\hat{\alpha}-\hat{\beta}_{[2:d]}^{T}\mu_{[2:d]}\right)\sqrt{\hat{\beta}^{T}\Sigma\hat{\beta}}$
$\Leftrightarrow$ $\tilde{y}\left(\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}-\sqrt{\hat{\beta}^{T}\Sigma\hat{\beta}}\right)>\left(\hat{\alpha}+\hat{\beta}_{[2:d]}^{T}\mu_{[2:d]}\right)\left(\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}-\sqrt{\hat{\beta}^{T}\Sigma\hat{\beta}}\right)+\left(\hat{\beta}_{1}\mu_{1}\right)\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}.$
Here, $\Phi$ is the standard normal cdf. If
$\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}-\sqrt{\hat{\beta}^{T}\Sigma\hat{\beta}}>0,$
group 0 is over-selected if and only if $C$ implies a threshold $\tilde{y}$ with
$\tilde{y}>\left(\hat{\alpha}+\hat{\beta}_{[2:d]}^{T}\mu_{[2:d]}\right)+\left(\hat{\beta}_{1}\mu_{1}\right)\frac{\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}}{\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}-\sqrt{\hat{\beta}^{T}\Sigma\hat{\beta}}}.$
If
$\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}-\sqrt{\hat{\beta}^{T}\Sigma\hat{\beta}}<0,$
group 0 is over-selected if and only if $C$ implies a threshold $\tilde{y}$ with
$\tilde{y}<\left(\hat{\alpha}+\hat{\beta}_{[2:d]}^{T}\mu_{[2:d]}\right)+\left(\hat{\beta}_{1}\mu_{1}\right)\frac{\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}}{\sqrt{\hat{\beta}_{[2:d]}^{T}\Sigma_{[2:d,2:d]}\hat{\beta}_{[2:d]}}-\sqrt{\hat{\beta}^{T}\Sigma\hat{\beta}}}.$
The proposition follows.
Corollary 2
We combine Proposition A.2 with the parameter estimates given in the proof of Proposition 2. For a high threshold $\tilde{y}$, the group with more under-reporting is over-selected if
$\mathds{V}\left[{\hat{\beta}_{[2:d]}^{T}Z_{[2:d]}}\right]>\mathds{V}\left[{\hat{\beta}Z}\right]$
$\Leftrightarrow$ $\sum_{i=2}^{d}\hat{\beta}_{i}^{2}\sigma_{i}^{2}+\sum_{i=2}^{d}\sum_{j=i+1}^{d}2\hat{\beta}_{i}\hat{\beta}_{j}\text{Cov}\left[{Z_{i}},{Z_{j}}\right]>\sum_{i=1}^{d}\hat{\beta}_{i}^{2}\sigma_{i}^{2}+\sum_{i=1}^{d}\sum_{j=i+1}^{d}2\hat{\beta}_{i}\hat{\beta}_{j}\text{Cov}\left[{Z_{i}},{Z_{j}}\right]$
$\Leftrightarrow$ $\hat{\beta}_{1}^{2}\sigma_{1}^{2}+\sum_{j=2}^{d}2\hat{\beta}_{1}\hat{\beta}_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<0.$
Recall that
$R^{2}=\frac{\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}$
and denote $S^{2}=\sum_{i=2}^{d}\rho(Z_{1},Z_{i})^{2}$. Then
$\frac{1}{1-R^{2}}=\frac{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]S^{2}}.$
Using Proposition 2 and inserting the parameter estimates gives
$\hat{\beta}_{1}^{2}\sigma_{1}^{2}+\sum_{j=2}^{d}2\hat{\beta}_{1}\hat{\beta}_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<0$
$\Leftrightarrow$ $\left(\frac{1}{1-R^{2}}\right)^{2}\left(\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}-\frac{R^{2}}{\mathds{E}\left[{\xi_{1}}\right]}\right)^{2}\beta_{1}^{2}\mathds{V}\left[{Z_{1}}\right]$
$+\frac{1}{1-R^{2}}\left(\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}-\frac{R^{2}}{\mathds{E}\left[{\xi_{1}}\right]}\right)\beta_{1}\sum_{j=2}^{d}2\left(\beta_{1}\frac{1}{1-R^{2}}\left(\frac{\text{Cov}\left[{Z_{1}},{Z_{j}}\right](\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2})}{\mathds{V}\left[{Z_{j}}\right]\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\right)+\beta_{j}\right)\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<0$
$\Leftrightarrow$ $\left(\frac{1}{1-R^{2}}\right)\left(\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}-\frac{R^{2}}{\mathds{E}\left[{\xi_{1}}\right]}\right)\beta_{1}^{2}\mathds{V}\left[{Z_{1}}\right]$
$+2\frac{1}{1-R^{2}}\beta_{1}^{2}\sum_{j=2}^{d}\left(\frac{\text{Cov}\left[{Z_{1}},{Z_{j}}\right](\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2})}{\mathds{V}\left[{Z_{j}}\right]\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\right)\text{Cov}\left[{Z_{1}},{Z_{j}}\right]+2\beta_{1}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<0$
$\Leftrightarrow$ $\frac{1}{1-R^{2}}\left(\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}-\frac{R^{2}}{\mathds{E}\left[{\xi_{1}}\right]}\right)\beta_{1}^{2}\mathds{V}\left[{Z_{1}}\right]$
$+2\beta_{1}^{2}\frac{1}{1-R^{2}}\frac{(\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2})\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]}\sum_{j=2}^{d}\left(\frac{\text{Cov}\left[{Z_{1}},{Z_{j}}\right]^{2}}{\mathds{V}\left[{Z_{j}}\right]\mathds{V}\left[{Z_{1}}\right]}\right)+2\beta_{1}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<0$
$\Leftrightarrow$ $\frac{\mathds{E}\left[{\xi_{1}}\right]\mathds{V}\left[{Z_{1}}\right]^{2}}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]S^{2}}(1-S^{2})\beta_{1}^{2}+2\beta_{1}^{2}\frac{(\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2})S^{2}\mathds{V}\left[{Z_{1}}\right]}{\mathds{V}\left[{Z_{1}\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}\mathds{V}\left[{Z_{1}}\right]S^{2}}+2\beta_{1}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<0$
$\Leftrightarrow$ $\beta_{1}^{2}\mathds{V}\left[{Z_{1}}\right]\frac{\mathds{V}\left[{Z_{1}}\right](1-S^{2})+2(1-\mathds{E}\left[{\xi}\right])\mathds{E}\left[{Z_{1}^{2}}\right]S^{2}}{\mathds{V}\left[{Z_{1}}\right](1-S^{2})\mathds{E}\left[{\xi_{1}}\right]+(1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]}+2\beta_{1}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<0$
where we used that $\mathds{V}\left[{Z_{1}\xi_{1}}\right]=\mathds{V}\left[{Z_{1}}\right]\mathds{E}\left[{\xi_{1}}\right]^{2}+\mathds{V}\left[{\xi_{1}}\right]\mathds{E}\left[{Z_{1}^{2}}\right]$ and $\mathds{V}\left[{\xi_{1}}\right]=\mathds{E}\left[{\xi_{1}}\right]-\mathds{E}\left[{\xi_{1}}\right]^{2}$. Note that the first term on the left side is always positive. Thus the inequality is fulfilled if and only if
$\text{sign}\left(\beta_{1}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]\right)=-1$
and
$2\beta_{1}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<-\beta_{1}^{2}\mathds{V}\left[{Z_{1}}\right]\frac{\mathds{V}\left[{Z_{1}}\right](1-S^{2})+2(1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]S^{2}}{\mathds{V}\left[{Z_{1}}\right](1-S^{2})\mathds{E}\left[{\xi_{1}}\right]+(1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]}$
$\Leftrightarrow$ $\frac{1}{\beta_{1}}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<-\frac{\mathds{V}\left[{Z_{1}}\right]^{2}(1-S^{2})+2(1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]\mathds{V}\left[{Z_{1}}\right]S^{2}}{2\mathds{V}\left[{Z_{1}}\right](1-S^{2})\mathds{E}\left[{\xi_{1}}\right]+2(1-\mathds{E}\left[{\xi_{1}}\right])\mathds{E}\left[{Z_{1}^{2}}\right]}$
$\Leftrightarrow$ $\frac{1}{\beta_{1}}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<-\frac{\mathds{V}\left[{Z_{1}}\right]^{2}(1-S^{2})+2(1-\mathds{E}\left[{\xi_{1}}\right])(\mathds{V}\left[{Z_{1}}\right]+\mathds{E}\left[{Z_{1}}\right]^{2})\mathds{V}\left[{Z_{1}}\right]S^{2}}{2\mathds{V}\left[{Z_{1}}\right](1-S^{2})\mathds{E}\left[{\xi_{1}}\right]+2(1-\mathds{E}\left[{\xi_{1}}\right])(\mathds{V}\left[{Z_{1}}\right]+\mathds{E}\left[{Z_{1}}\right]^{2})}.$
Since we know that the fraction on the right side is always positive, this can be rewritten as presented in the corollary. If the inequality is not fulfilled, the group with more under-reporting is under-selected at a high threshold.
Lemma 1
In the setting of the Lemma, we can write
$\mathds{E}_{\xi}\left[l(f,X,y)\right]$ $=\mathds{E}_{\xi}\left[\frac{1}{m}l(f,X,y)-\frac{1-m}{m}l(f,[0,X_{[2:d]}]^{T},y)\right]$
$=\frac{1}{m}\mathds{E}_{\xi}\left[l(f,X,y)\right]-\frac{1-m}{m}l(f,[0,z_{[2:d]}]^{T},y)$
$=\frac{1}{m}\left(P(\xi_{1}=1)l(f,z,y)+P(\xi_{1}=0)l(f,[0,z_{[2:d]}]^{T},y)\right)-\frac{1-m}{m}l(f,[0,z_{[2:d]}]^{T},y)$
$=l(f,z,y)+\frac{1-m}{m}l(f,[0,z_{[2:d]}]^{T},y)-\frac{1-m}{m}l(f,[0,z_{[2:d]}]^{T},y)$
$=l(f,z,y).$
Here, the first equality holds since only the first feature has under-reporting and the second equality holds because $Z\bot G$ which implies $Z\bot\xi$.
Lemma A.3
Follows the same as Lemma 1. Instead of under-reporting completely at random, the under-reporting is completely at random within group $g$.
Lemma 2
Since $f$ is linear and under-repoting only occurs in the first feature, the expected prediction error for an imputation value $x_{1}^{\prime}$ can be written as
$R(f)$ $=\mathds{E}_{X}\left[(f(X^{\prime})-Y)^{2}\right]$
$=\beta_{1}^{2}\mathds{E}_{X}\left[(X_{1}^{\prime}-Z_{1})^{2}\right]$
$=\beta_{1}^{2}P(X_{1}=0)\mathds{E}_{X}\left[(X_{1}^{\prime}-Z_{1})^{2}\mid X_{1}=0\right]+\beta_{1}^{2}P(X_{1}\neq 0)\mathds{E}_{X}\left[(X_{1}^{\prime}-Z_{1})^{2}\mid X_{1}\neq 0\right]$
$=\beta_{1}^{2}P(X_{1}=0)\mathds{E}_{Z}\left[(x_{1}^{\prime}-Z_{1})^{2}\mid X_{1}=0\right]+\beta_{1}^{2}P(X_{1}\neq 0)\mathds{E}_{Z}\left[(Z_{1}-Z_{1})^{2}\mid X_{1}\neq 0\right]$
$=\beta_{1}^{2}P(X_{1}=0)\left(x_{1}^{\prime 2}-2x_{1}^{\prime}\mathds{E}\left[Z_{1}\mid X_{1}=0\right]+\mathds{E}\left[Z_{1}^{2}\mid X_{1}=0\right]\right).$
Then
$\frac{dR(f)}{dx_{1}^{\prime}}=\beta_{1}^{2}P(X_{1}=0)(2x_{1}^{\prime}-2\mathds{E}\left[{Z_{1}\mid X_{1}=0}\right])\overset{!}{=}0$
$\Leftrightarrow$ $x_{1}^{\prime}=\mathds{E}\left[{Z_{1}\mid X_{1}=0}\right].$
We implicitly assume that $\beta_{1}\neq 0$ and the probability of 0-entries is positive.
Lemma A.4
Recall that $G\sim\text{Bern}(r)$. Similar to the proof of Lemma 2, the expected prediction error can be written as
$R(f)$ $=\mathds{E}_{X}\left[(f(X^{\prime})-Y)^{2}\right]$
$=\beta_{1}^{2}P(X_{1}=0)\mathds{E}_{Z}\left[(X_{1}^{\prime}-Z_{1})^{2}\mid X_{1}=0\right]$
$=\beta_{1}^{2}P(X_{1}=0)\left(r\mathds{E}_{Z}\left[(X_{1}^{\prime}-Z_{1})^{2}\mid X_{1}=0,G=1\right]+(1-r)\mathds{E}_{Z}\left[(X_{1}^{\prime}-Z_{1})^{2}\mid X_{1}=0,G=0\right]\right)$
$=\beta_{1}^{2}P(X_{1}=0)\left(r\mathds{E}_{Z}\left[(x_{1}^{\prime 1}-Z_{1})^{2}\mid X_{1}=0,G=1\right]+(1-r)\mathds{E}_{Z}\left[(x_{1}^{\prime 0}-Z_{1})^{2}\mid X_{1}=0,G=0\right]\right).$
This prediction error is minimal when
$\frac{dR(f)}{dx_{1}^{\prime g}}=\beta_{1}^{2}P(X_{1}=0)P(G=g)(2x_{1}^{\prime g}-2\mathds{E}\left[{Z_{1}\mid X_{1}=0,G=g}\right])\overset{!}{=}0$
$\Leftrightarrow$ $x_{1}^{\prime g}=\mathds{E}\left[{Z_{1}\mid X_{1}=0,G=g}\right]$
for $g\in\{0,1\}$.
Appendix D Supplementary Figures
Name #Obs. #Feat. Groups Binary outcomes
COMPAS (Angwin et al., 2016)
7,214
6
Race (51% African-American, 49% other), Gender (81% male, 19% female)
Two-year recidivism, violent recidivism
German credit (Repository, 1994)
1,000
19
Gender (69% male, 31% female)
Good credit
ACS Income (CA, 2018) (Ding et al., 2021)
195,665
6
Race (62% White, 38% other), Gender (53% male, 47% female)
Yearly income over \$50,000
Birth data
39,365
51
Medicaid (no 72%, yes 28%), Race (African-American 21%, other 79%)
Child placed in foster care within 3 years
Table D.1. Statistics of the datasets used in experiments. Data is split randomly into 80% for training and 20% for testing. For the first three datasets, we iterate over all outcome types, groups, and numerical features for missingness injection.
Figure D.1. Excess selection rate of racial group Other (i.e. not African-American) at different selection rates of the whole population with synthetic two-year recidivism outcomes using the COMPAS dataset. Each panel represents a feature that has been corrupted by under-reporting in independent runs of the experiment. Feature under-reporting is added to the Other group with 0-90% missing in 10 percentage point increments. The black curves show performance when excluding the whole feature column from modeling. Results are reported as averages over 30 runs on a test set. Shaded areas correspond to one standard deviation in each direction of the mean.
[…]
Figure D.2. Test set $R^{2}$ over varying levels of feature under-reporting in the ACS Income dataset. Results are reported as averages over 50 runs on the test set. Variation in results was minimal.
Figure D.3. Parameter estimates over varying levels of feature under-reporting in the ACS Income dataset. Each panel indicates a different feature selected for under-reporting injection. Points indicate the true parameters from the semi-synthetic ground truth model. Results are reported as averages over 50 runs. Variation in estimates over runs was minimal. Note that only estimates for continuous features are displayed and the estimated parameters for the levels of the categorical variables worker class, marital status, and relationship to reference person are omitted for readability.
Figure D.4. Multiple imputation excess selection rate of racial group Other (i.e. not African-American) at different selection rates of the whole population with synthetic two-year recidivism outcomes using the COMPAS dataset. Results are reported as averages over 30 runs on a test set. Shaded areas correspond to one standard deviation in each direction of the mean.
Figure D.5. Excess selection rate of racial group Other (i.e. not African-American) when training on rows without 0-entries at different selection rates of the whole population with synthetic two-year recidivism outcomes using the COMPAS dataset. Results are reported as averages over 30 runs on a test set. Shaded areas correspond to one standard deviation in each direction of the mean.
Figure D.6. Test set $R^{2}$ of different solution approaches using the COMPAS dataset with synthetic two-year recidivism outcomes. Under-reporting is injected into the feature ‘priors count’ of group Other (i.e. not African-American). Results are reported as averages over 30 runs. Shaded areas correspond to one standard deviation.
Appendix E Under-reporting rate estimation
We estimate under-reporting rates using the procedure described in Section 6 where $V$ is taken to be the 80% training data fold conditioned on the group with under-reporting. Half of the data is used for training of $h$ while the other half is used as evaluation data to compute $\hat{m}_{g}$. As model class for $h$, we use XGBoost classifiers with 100 trees of maximum depth 3, and learning rate 0.1. Figure E.7 depicts the results of the estimation procedure for under-reporting in the feature priors count for the racial group Other and synthetic two-year recidivism outcomes using the COMPAS dataset. We see that estimation of $m_{g}$ works particularly well when under-reporting in the feature is high. When the feature is fully observed, i.e. $m_{g}=1$, the estimator returns $\hat{m}_{g}=0.811$ on average. We hypothesize that this is due to the high share of true 0-entries occurring in the priors count feature which impacts the estimator more when fewer values are missing due to under-reporting.
Figure E.7. True vs. estimated observed rate (i.e. 1 - under-reporting rate) of feature priors count in racial group Other (i.e. not African-American) with sythetic two-year recidivism outcomes using the COMPAS dataset. Results are reported as averages over 30 runs. Shaded areas correspond to one standard deviation in each direction of the mean. The black line shows $y=x$ for comparison.
Appendix F Additional experiments and results
F.1. German credit data
Figure F.8. Excess selection rate of male group at different selection rates of the whole population with synthetic outcomes using the German credit dataset. Each panel represents a feature that has been corrupted by under-reporting in independent runs of the experiment. Feature under-reporting is added to the male group with 0-90% missing in 10 percentage point increments. The black curves show performance when excluding the whole feature column from modeling. Results are reported as averages over 50 runs on a test set. Shaded areas correspond to one standard deviation in each direction of the mean. Note that feature under-reporting is only injected into continuous features and models are estimated using the displayed features as well as the available categorical features checking account status, credit history, purpose, savings, employment, marital status, type of owned property, other installment plans, housing type, and job type.
Our experiments suggest that addition of feature under-reporting to one of the two gender groups in the German credit dataset has only marginal fairness implications. Figure F.8 depicts the results for synthetic outcomes and addition of different amounts of under-reporting to the features of the male group. We can see that, for any of the considered features, the amount of under-reporting injected has little to no effect on the excess selection rate of the male group. However, when selecting rates of around 10-15% from the whole population any amount of under-reporting in the installment feature appears to results in a slight over-selection of the male group. The installment feature in the German credit dataset is discretized into four values with lower values indicating a higher installment rate. Incorrectly observed 0-values may thus suggest a high installment rate which is indicative of good credit.
F.2. Beyond the noise-free setting
Motivation
The experiments on the publicly available datasets discussed in Sections 7 and 8 rely on semi-synthetic outcomes that are computed as deterministic linear functions of correctly measured features. The implicit simplifying assumptions are that, without feature under-reporting, a linear model on the data can retrieve the true data generating model $f(X)=\alpha+\beta^{T}Z$ and the exact outcomes $Y$ as recorded in the data. This modeling choice facilitates isolation of the effect of feature under-reporting by explicitly excluding potential effects of model misspecification and regression noise. In real-life applications, we can generally not predict outcomes exactly even if correctly measured features are available. In the following additional set of experiments, we loosen the assumption of a noise-free regression setting to allow for more general settings.
Experimental setup
We follow a similar experimental setup as described in Section 7.1. Instead of relying on noise-free regression labels $Y=\alpha+\beta^{T}Z$, we add some noise back into the system by setting
$Y=\alpha+\beta^{T}Z+\varepsilon.$
Here, $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ is i.i.d. and assumed to have mean zero. Like before, fitting a linear regression of $Y$ on features $Z$ with sufficient data yields the correct parameter estimates $\hat{\beta}=\beta$ and $\hat{\alpha}=\alpha$. However, in contrast to the noise-free setting, the prediction model $\hat{Y}_{Z}=\alpha+\beta^{T}Z$ can only retrieve labels $Y$ up to random noise. The $R^{2}$ of this prediction model can be controlled via the variance $\sigma^{2}$ by setting
$\sigma^{2}=\frac{1-R^{2}}{R^{2}}\mathds{E}\left[{\left((\alpha+\beta^{T}Z)-\mathds{E}\left[{\alpha+\beta^{T}Z}\right]\right)^{2}}\right].$
We experiment with $R^{2}$ values between 0.1 and 1.0 in increments of 0.1. Instead of comparing predictions $\hat{Y}_{Z}$ to predictions under feature under-reporting $\hat{Y}_{X}$, we compare thresholded versions of outcomes $Y$ and $\hat{Y}_{X}$ directly to measure both the impact of under-reporting and regression noise.
[…]
Figure F.9. Excess selection rate under under-reporting over true labels $Y$ with different levels of $R^{2}$ for the model on correctly measured features $Z$. Low $R^{2}$ indicates a high level of noise and vice versa. Under-reporting injected into the features of group Other (i.e. not African-American) in the COMPAS dataset. Results are reported as averages over 30 simulation runs with shaded areas representing one standard deviation in each direction.
Results
Figure F.9 depicts a subset of the results for the COMPAS dataset. Comparing against the results of the noise-free setting summarized in Figure D.1, we observe that the group Other is under-selected to a greater extend with additional noise. Under-selection occurs even if no feature under-reporting is added (dark blue curves) and increases with increasing noise, i.e. decreasing $R^{2}$ of the model on $Z$. On a high level, this occurs because the predictions $\hat{Y}_{Z}$ concentrate more closely around their group-level means as compared to the true values $Y$. The mean of $\hat{Y}_{Z}$ is smaller for the group Other than the group African-American which leads to under-selection of the group Other as compared to the true $Y$ at many thresholds. We note that the group-level variances in outcomes $Y$ and predictions $\hat{Y}_{Z}$ play a role in this dynamic as well. The isolated effect of feature under-reporting in the studied setting appears to be similar to the effect in the noise-free setting. As under-reporting is introduced into the group Other via the feature ‘priors count’, the group Other is further under-selected. The more under-reporting is injected, the more the group is under-selected. The magnitude of under-selection due to under-reporting is comparable across different levels of regression noise. Overall, the results give us some insight into what to expect in more realistic settings of feature under-reporting. Instead of selection rate disparities that are exclusively due to differential feature under-reporting, disparities in the studied setting also depend on regression noise which, together, leads to increased disparities overall.
F.3. Possibility of decreasing disparities
We conduct our main experiments on three publicly available datasets, i.e. COMPAS data (Angwin et al., 2016), German credit data (Repository, 1994), and ACS Income data (Ding et al., 2021), where each numerical feature is considered for the effect of under-reporting. As discussed in Section 8, the results suggest that, if an effect is present, feature under-reporting generally leads to under-selection of the group with under-reporting which aligns with Case 2 from the theoretical derivations in Section 5. If the group with under-reporting aligns with the group that is less frequently selected in the ground truth model, this implies that differential feature under-reporting leads to increased selection rate disparities.
All three datasets have a numerical age feature which was considered for under-reporting but omitted for the discussion of results in the main text. In contrast to most other features (e.g. the counts in the COMPAS data), the default value of 0 is somewhat unintuitive for age and lies outside of the feature’s support in each of the datasets. Studying the effect of fitting a model on differentially available data directly is less compelling in this setting since we essentially have indicators for under-reporting and could hope to use missing data methods like imputation directly. Nevertheless, we discuss the results for feature under-reporting in age for the COMPAS dataset in the following as it presents the only empirical example for decreasing disparities we encounter in our experiments.
Figure F.10 depicts the parameter estimates and excess selection rate of group Other (i.e. not African-American) when fitting a model on data with feature under-reporting in the feature ‘age’ for group Other. We see that under-reporting in this setting leads to over-selection of the group with under-reporting. This over-selection is increasing with increasing levels of under-reporting. As the figure shows, the regression parameter for age is negative with an attenuation effect when under-reporting is injected. This means that in the semi-synthetic ground truth model and in the prediction models under under-reporting younger defendants are more likely to reoffend than older defendants. The feature correlations between age and juvenile crime counts (felony, misdemeanor, and other) are negative in the data while the correlation between age and the feature ‘priors count’ is positive. This leads to parameter estimates that are increasing for increasing under-reporting in age for juvenile crime counts and decreasing for increasing under-reporting for priors count exactly as predicted by the theoretical analyses in Proposition 3. Ultimately, this example shows how, in some settings, disparities may decrease as a function of under-reporting which aligns with Case 1 from the theoretical discussions in Section 5. However, the example presented here is somewhat artificial and we find that typically disparities are increasing with differential feature under-reporting.
Figure F.10. Excess selection rate of group Other (i.e. not African-American) at different population selection rates with synthetic outcomes using the COMPAS dataset, and the respective parameter estimates. Under-reporting is added to the feature ‘age’ in group Other. Results are reported as averages over 30 simulation runs with shaded areas representing one standard deviation in each direction. Note that parameter estimates are only displayed for continuous count features and age to preserve readability. The models additionally take sex and the categorical feature charge degree into consideration.